### v2

# ReDDiT: Rehashing Noise for Discrete Visual Generation

Abstract
Discrete diffusion models are gaining traction in the visual generative area for their efficiency and compatibility. However, the pioneered attempts still fall behind the continuous counterparts, which we attribute to the noise (absorbing state) design and sampling heuristics. In this study, we propose the rehashing noise framework for discrete diffusion transformer (termed ReDDiT), to extend absorbing states and improve expressive capacity of discrete diffusion models. ReDDiT enriches the potential paths that latent variables traverse during training with randomized multi-index corruption. The derived rehash sampler, which reverses the randomized absorbing paths, guarantees the diversity and low discrepancy of the generation process. These reformulations lead to more consistent and competitive generation quality, mitigating the need for heavily tuned randomness. Experiments show that ReDDiT significantly outperforms the baseline (reducing gFID from 6.18 to 1.61) and is on par with the continuous counterparts with higher efficiency.
1 University of Chinese Academy of Sciences 2 China Mobile Research Institute
matianren18@mails.ucas.ac.cn   qxye@ucas.ac.cn
Project Page: https://github.com/martian422/ReDDiT
1 Introduction
Diffusion has been a highly competitive approach for generative workloads [8, 30, 21], offering strong bidirectional perception and well-structured mechanisms [48] for global control over content. In the continuous domain, diffusion transformers (DiTs) [28], which progressively refine image latents from Gaussian noise, have achieved impressive and scalable results. Despite the popularity of DiTs, the community shows a growing interest in discrete diffusion models in recent years [16, 40]. This trend is based on their practical advantages, e.g., compatibility with language models for the indexable codebooks and efficiency for the ability to generate multiple tokens at each inference. Early endeavors [4, 5, 12] mainly pursue efficiency through integrating visual tokenizers and BERT-style [mask] tokens [7]. Recent studies [2, 45] improved the generation quality, demonstrating great potential of visual generation in discrete domain.
Despite the progress, discrete diffusion methods remain lagging behind their continuous counterparts. State-of-the-art approaches, e.g., masked visual token models (MVTMs) [45], is puzzeld by its mask design and predict-remask sampler, which restricts the model’s expressive capacity and makes the prediction performance sensitive to adaptions even with extensive training, Fig. 1(upper). Moreover, when paired with large-vocabulary codebooks from high-fidelity modern tokenizers, these models encounter challenges such as slower sampling speeds and numerical inaccuracy [49].
Figure 1: Comparison of ReDDiT with the baseline discrete model (MVTM). MVTMs rely on score-based remasking strategies with Gumbel-max to sample from logits, which leads to lower token diversity and suboptimal token selection. In contrast, ReDDiT introduces a principled, low-discrepancy rehashing mechanism that leverages softmax-based probabilities, enabling diverse, high-quality sampling through a learned distribution. (Best viewed in color)
To address these limitations, we propose two key hypotheses. First, while previous methods learn to recover plausible tokens from a monotonous [mask] canvas, this noise design may not be well-suited for discrete visual generation. In continuous diffusion, Gaussian noise is used to progressively degrade the input to learn a smooth distribution shift  [15, 22]. Discrete masking mimics this paradigm by collapsing all masked tokens into a single absorbing state, which however, lacks the variability of Gaussian noise, in terms of both vocabulary richness and latent diversity. Consequently, the discrete process offers a far coarser signal, which limits its ability to capture and represent diverse data distributions [33, 1]. Moreover, while continuous diffusion models introduce stochasticity at every inference step through noise injection, discrete unmasking is inherently binary: tokens are either masked or deterministically decoded. This rigid mechanism further constrains the flexibility of sample refinement during generation.
Second, the predict-remask sampler of MVTMs introduces a form of handcrafted randomness, which is often implemented using Gumbel-max, to approximate sampling diversity. However, this sampler may compromise the probabilistic fidelity of the generation, and the need to carefully balance the token numbers decoded per step to mitigate accumulation errors leads to redundant sampling passes. As a result, the Gumbel-max has evolved to a heavily tuned time variant trick, leading to unstable performance, particularly when scaled to large-vocabulary codebooks. The above factors, rather than quantization alone, attribute to the performance gap between discrete and continous models.
In this study, we propose a discrete diffusion model with an elaborate rehashing noise design, Fig. 1(lower). Our approach, termed ReDDiT, addresses the limitations of the uni-mask design by redefining absorbing states towards larger representational capacity, through enriching the potential paths that latent variables can traverse during diffusion. Specifically, we expand the masks to multiple indices along with the codebook and randomize them during data corruption. A rehash sampler is also derived with principled discrete diffusion theories to reverse the diffusion path for generation, guaranteeing the diversity and low discrepancy of the sampling process. We demonstrate that this rehashed noise helps learning a superior and flatter lower error bound while eliminating reliance to hyper-parameterized randomness during sampling.
We further revisit the commonly used discrete diffusion objective (e.g., the masked cross-entropy loss) and update it with empirical modifications. By adopting an improved ELBO [31, 36] with representation alignment (RepA) [46] loss, we achieve better training efficiency and generation quality. Moreover, ReDDiT aligns with recent advances in discrete flow matching [11, 34], enabling token refreshment during sampling without training post-correction models [19].
2 Related Work
Diffusion Models.
Diffusion models [15, 37] have emerged as a powerful class of generative methods that learn data distributions by reversing a gradual noising process over time. These models are primarily designed for continuous domains such as images [8, 10, 28], defining a forward process that transforms data $x_{0}$ into noise $x_{1}$: $x_{t}\sim\mathcal{N}(\sqrt{\alpha_{t}}x_{0};(1-\alpha_{t})\mathbf{I})$ where $\alpha_{t}$ controls the noise schedule. The generative (reverse) process learns a denoising model $p_{\theta}(x_{s}\mid x_{t})$, often parameterized via a neural network $\theta$ to predict either noise or clean data.
Discrete Diffusion Models.
Discrete diffusion has been previously governed by masked visual token models (MVTMs) [4, 5, 12, 44, 45]. This model leverages a BERT-style [mask] token to corrupt the tokenized image sequence and trained the network with a simple cross-entropy loss on masked tokens, resulting in a score-based prediction. It generates tokens in a non-autoregressive fashion, by remasking the tokens with least scores at each inference as depicted in Alg. 1.
Recent studies unlocked the principled discrete diffusion model (DDM) [31, 36] and discrete flow-matching (DFM) [11, 34], which adapt the Markov chain theory, enabling generation over text [26, 23], molecules[34], and other discrete representations [1, 24]. Unlike MVTMs, the principled DDM and DFM mostly derive a time-weighted cross-entropy loss to supervise the training procedure and apply a gradual unmasking method based on probabilities.
3 Methodology
For self-containment, we first review the DDM theory in Sec. 3.1. We reformulate its diffusion dynamics and introduce rehashing noise for ReDDiT in Sec. 3.2. Connection and comparison of ours and other discrete diffusion models are discussed in Sec. 3.3.
3.1 Preliminary: Discrete Diffusion Model
DDM defines a forward process over discrete variables by gradually corrupting the image tokens into absorbing states (masks) through a continuous-time Markov process. Assume that the data consists of tokens from a finite vocabulary $\mathcal{V}$, $x\in\mathcal{V}^{L}$ represents a sequence of tokens ($e.g.$, an image tokenized into indices) with length $L$. We denote the clean data as $x_{t=0}$ ($x_{0}$ for short), and noise it gradually as $t\rightarrow 1$. DDM defines an absorbing token $\mathbf{m}\in\mathcal{V}$, such that once a token is noised to $\mathbf{m}$ it remains unchanged. At the terminal time $t=1$, $x_{t}$ fully transits to $\mathbf{m}^{L}$, which means $x_{1}^{i=1\sim L}=\mathbf{m}$.
Let $\alpha_{t}$ be the noise scheduler (a monotonically decreasing survival function that satisfies $\alpha_{0}=1,\alpha_{1}=0$ ). For $0\leq s<t\leq 1$, denote $\alpha_{t|s}^{\leftarrow}=\alpha_{t}/\alpha_{s}$, the forward corruption process is governed by a continuous-time transition kernel $q(x_{t}^{i}\mid x_{s}^{i})$ on the $i$-th element, as
$q(x_{t}^{i}\mid x_{s}^{i})=\begin{cases}1-\alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=\mathbf{m},x_{s}^{i}\neq\mathbf{m},\\ \alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}\neq\mathbf{m},\\ 1,&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}=\mathbf{m},\\ 0,&\text{otherwise}.\end{cases}$ (1)
Given the transition kernel, the corrupted data distribution at time $t$ is determined  [31] as
$x_{t}\sim\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{m}^{L}$ (2)
The generative model learns the reverse process $p_{\theta}(x_{s}\mid x_{t})$, which denoises sample $x_{t}$ at arbitrary time $t\in(0,1]$ to a less noised state $x_{s}$ at time $s<t$. Denote $\delta(x_{t}^{i},\textbf{m})$ as the indicator function that only computes on masked tokens, and $\alpha_{t}^{\prime}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$, the learning objective is derived  [36, 31] as
$\mathcal{L}_{\text{DDM}}=-\mathbb{E}_{x_{0},\ x_{t}}\int_{t=0}^{t=1}[\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t})]\mathrm{d}t\ .$ (3)
For a linear scheduler, Eq. 3 is simplified via variable substitution [36, 23] to an equivalent form:
$\mathcal{L}_{\text{DDM-linear}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}[\frac{1}{t}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t})]\ .$ (4)
For conditional generation, class information $c$ (e.g., labels or text prompts) is introduced to the denoising model as additional input. Following classifier-free guidance [14], the model is trained with a random drop of labels, and the prediction is interpolated at inference:
$\hat{p}_{\theta}(x_{t},c)=\ p_{\theta}(x_{t},\varnothing)+w\cdot(\,p_{\theta}(x_{t},c)-\ p_{\theta}(x_{t},\varnothing)),$ (5)
where $w\geq 0$ controls the guidance strength.
Figure 2: Visualization of pixel and latent spaces. $m$ denotes the number of enriched noise indices. 3D t-SNE plot (right) is used solely to separate valid and noise tokens for clustering illustration.
3.2 Discrete Diffusion with Rehashing Noise
The ordinal structure inherent in discrete data provides a valuable inductive bias for designing transition kernels in diffusion dynamics. Prior studies [1, 3] show that assigning higher transition probabilities to neighboring pixel values—forming a discrete Gaussian-like noise—outperforms the single absorbing state approach on pixel-level datasets like CIFAR-10. However, when using visual tokenizers, the structure of discretized latents is learned rather than pre-defined, making such ordinal assumptions inapplicable. This insight motivates us to extend the conventional mask token into a set of indices, and reverse the diffusion path with noise rehashing. Such a procedure allows the model to optimize its embedding space during training, enhancing its ability to model flexible and data-driven noise structures. We visualize the learned distributions in Fig. 2 (right).
Reformulation.
Given $d$ categories, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be its one-hot vector where the $i$-th value is $1$. We denote =$\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution (known as $d$-simplex), and a basis for absorbing states with capacity $m$: =$\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$. The sum of $\mathcal{E}$ and $\mathcal{M}$ can be denoted as
$\mathcal{V}_{(d,m)}\triangleq\left\{\mathbf{v}_{(i,j)}\in\mathbb{R}^{d+m}\,\middle|\,\mathbf{v}_{(i,j)}=\begin{cases}\mathbf{e}_{i}\oplus\mathbf{0}_{m},&\text{for }i=1,\ldots,d,\ j=0\\ \mathbf{0}_{d}\oplus\mathbf{m}_{j},&\text{for }j=1,\ldots,m,\ i=0\end{cases}\right\}.$ (6)
We further denote the subspace $\mathcal{E}_{d},\ \mathcal{M}_{m}\in\mathcal{V}_{(d,m)}$ which contain either valid or mask tokens, as
$\mathcal{E}_{d}=\left\{\mathbf{v}_{(i,0)}\in\mathcal{V}_{(d,m)}\,\middle|\,i=1,\ldots,d\right\},\ \mathcal{M}_{m}=\left\{\mathbf{v}_{(0,j)}\in\mathcal{V}_{(d,m)}\,\middle|\,j=1,\ldots,m\right\}.$ (7)
To exploit visits across all the possible paths, we rewrite the transition kernel defined by Eq. 1 into
$q(x_{t}^{i}\mid x_{s}^{i})=\begin{cases}1-\alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=x_{s}^{i},\ x_{s}^{i}\notin\mathcal{M}_{m},\\ 1/m,&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise}.\end{cases}$ (8)
We introduce the essential transition matrix to model the process. Let $\mathbf{I}_{(d,m)}$, $\mathbf{M}_{(d,m)}$ and $\bm{\pi}_{(d,m)}$ be matrices in $\mathbb{R}^{(d+m)\times(d+m)}$, defined as
$\mathbf{I}_{(d,m)}=\begin{bmatrix}I_{d}&0\\ 0&0\end{bmatrix},\quad\mathbf{M}_{(d,m)}=\begin{bmatrix}0&\frac{1}{m}\mathbf{1}_{d\times m}\\ 0&0\end{bmatrix},\quad\bm{\pi}_{(d,m)}=\begin{bmatrix}0&0\\ 0&\frac{1}{m}\mathbf{1}_{m}\mathbf{1}_{m}^{\top}\end{bmatrix}$ (9)
where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix, and $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of ones. The transition matrix $Q_{t|s}\in\mathbb{R}^{(d+m)\times(d+m)}$ corresponding to Eq. 8 is
$Q_{t|s}=\alpha_{t|s}^{\leftarrow}\mathbf{I}_{(d,m)}+(1-\alpha_{t|s}^{\leftarrow})\mathbf{M}_{(d,m)}+\bm{\pi}_{(d,m)}$ (10)
which can be demonstrated intuitively:
$Q_{t|s}=\begin{bmatrix}\alpha_{t|s}^{\leftarrow}&0&\cdots&0&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}&\cdots&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}\\ 0&\alpha_{t|s}^{\leftarrow}&\cdots&0&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}&\cdots&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\alpha_{t|s}^{\leftarrow}&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}&\cdots&\frac{1-\alpha_{t|s}^{\leftarrow}}{m}\\ 0&0&\cdots&0&\frac{1}{m}&\frac{1}{m}&\cdots&\frac{1}{m}\\ 0&0&\cdots&0&\frac{1}{m}&\frac{1}{m}&\cdots&\frac{1}{m}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0&\frac{1}{m}&\frac{1}{m}&\cdots&\frac{1}{m}\end{bmatrix}$
[…]
The corrupted data distribution is a direct derivative of Eq. 10 by setting $s=0$:
$x_{t}$ $=x_{0}Q_{t|0}$
$=\alpha_{t}x_{0}\mathbf{I}_{(d,m)}+(1-\alpha_{t})x_{0}\mathbf{M}_{(d,m)}+x_{0}\bm{\pi}_{(d,m)}$
$=\alpha_{t}x_{0}+(1-\alpha_{t})x_{0}\mathbf{M}_{(d,m)}$
$\sim\alpha_{t}x_{0}+(1-\alpha_{t})\,\text{U}(\mathcal{M}_{m}^{L})$ (11)
where $\text{U}(\mathcal{M}_{m}^{L})$ is the uniform distribution on $\mathcal{M}_{m}^{L}$.
Rehash Sampling.
To generate a sequence of length $L$, the reverse process starts with $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$. Let $\mathbf{a}\odot\mathbf{b}$ denote the Hadamard product between two vectors $\mathbf{a}$ and $\mathbf{b}$, the reverse process is inferred from D3PM deduction [1] as:
[…]
$=\dfrac{[\alpha_{t|s}^{\leftarrow}\mathbf{I}_{(d,m)}x_{t}+(1-\alpha_{t|s}^{\leftarrow})\mathbf{M}_{(d,m)}x_{t}+\bm{\pi}_{(d,m)}x_{t}]\odot[\alpha_{s}x_{0}+(1-\alpha_{s})\mathbf{M}_{(d,m)}^{\top}x_{0}]}{x_{t}^{\top}[\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{M}_{(d,m)}^{\top}x_{0}+\pi_{(d,m)}^{\top}x_{0}]}$
$=\dfrac{[\alpha_{t|s}^{\leftarrow}\mathbf{I}_{(d,m)}x_{t}+(1-\alpha_{t|s}^{\leftarrow})\mathbf{M}_{(d,m)}x_{t}+\bm{\pi}_{(d,m)}x_{t}]\odot[\alpha_{s}x_{0}+(1-\alpha_{s})\mathbf{M}_{(d,m)}^{\top}x_{0}]}{\alpha_{t}x_{t}^{\top}x_{0}+(1-\alpha_{t})x_{t}^{\top}\mathbf{M}_{(d,m)}^{\top}x_{0}}$ (12)
We consider the separate cases: $x_{t}^{i}=x_{0}^{i}$ and $x_{t}^{i}\in\mathcal{M}_{m}$.
Case 1.
For $x_{t}^{i}=x_{0}^{i}$, Eq. 3.2 is simplified as
$q(x_{s}^{i}\mid x_{t}^{i}=x_{0}^{i})$ $=\dfrac{\alpha_{t|s}^{\leftarrow}x_{0}^{i}\odot\alpha_{s}x_{0}^{i}}{\alpha_{t}x_{0}^{i\ \top}x_{0}^{i}}$
$=1$ (13)
Case 2.
For $x_{t}^{i}\in\mathcal{M}_{m}$, we have
$q(x_{s}^{i}\mid x_{t}^{i}\in\mathcal{M}_{m})$ $=\dfrac{[(1-\alpha_{t|s}^{\leftarrow})\mathbf{M}_{(d,m)}x_{t}^{i}+\bm{\pi}_{(d,m)}x_{t}^{i}]\odot[\alpha_{s}x_{0}+(1-\alpha_{s})\mathbf{M}_{(d,m)}^{\top}x_{0}]}{(1-\alpha_{t})x_{t}^{i\ \top}\mathbf{M}_{(d,m)}^{\top}x_{0}}$
$=\dfrac{[(1-\alpha_{t|s}^{\leftarrow})\alpha_{s}\mathbf{M}_{(d,m)}x_{t}^{i}\odot x_{0}+\bm{\pi}_{(d,m)}(1-\alpha_{s})x_{t}^{i}\odot\mathbf{M}_{(d,m)}^{\top}x_{0}]}{(1-\alpha_{t})x_{t}^{i\ \top}\mathbf{M}_{(d,m)}^{\top}x_{0}}$
$=\dfrac{(\alpha_{s}-\alpha_{t})\mathbf{M}_{(d,m)}x_{t}^{i}\odot x_{0}+(1-\alpha_{s})\bm{\pi}_{(d,m)}x_{t}^{i}\odot\mathbf{M}_{(d,m)}^{\top}x_{0}}{(1-\alpha_{t})x_{t}^{i\ \top}\mathbf{M}_{(d,m)}^{\top}x_{0}}$ (14)
Denote $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$ to describe the reverse transition rate, and we have
$q(x_{s}^{i}\in\mathcal{M}_{m}\mid x_{t}^{i}\in\mathcal{M}_{m})=\dfrac{1-\alpha_{s}}{m(1-\alpha_{t})}=\dfrac{\alpha_{t|s}^{\rightarrow}}{m}$ (15)
$q(x_{s}^{i}\notin\mathcal{M}_{m}\mid x_{t}^{i}\in\mathcal{M}_{m})=\dfrac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}=1-\alpha_{t|s}^{\rightarrow}$ (16)
Combining case 1 with case 2, we have
$q(x_{s}^{i}\mid x_{t}^{i})=\begin{cases}1,&\text{if }x_{s}^{i}=x_{t}^{i},\ x_{t}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\rightarrow}/m,&\text{if }x_{s}^{i}\in\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 1-\alpha_{t|s}^{\rightarrow},&\text{if }x_{s}^{i}\notin\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise.}\\ \end{cases}$ (17)
Following MDLM’s deduction, assume that the denoising network can reconstruct $x_{0}$ perfectly, we use $p_{\theta}(x_{t})$ to approximate this reverse process for complex sequences, and get
$p_{\theta}^{i}(x_{s}^{i}|x_{t})=\begin{cases}1,&\text{if }x_{s}^{i}=x_{t}^{i},\ x_{t}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\rightarrow}/m,&\text{if }x_{s}^{i}\in\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ (1-\alpha_{t|s}^{\rightarrow})p_{\theta}^{i}(x_{t}),&\text{if }x_{s}^{i}\notin\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise.}\\ \end{cases}$ (18)
The subsequent latents $x_{t}$ are generated by discretizing the reverse timeline $T$ into $K$ steps. We denote this schedule as $T^{1:K+1}$ such that $T^{1}=1$ and $T^{K+1}=\varepsilon$, with $\varepsilon$ being an arbitrarily small positive constant. Inspired by MD4 [36], we use non-uniform time intervals for efficient sampling.
Algorithm 1 MVTM Sampling
1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$,
2:Settings: codebook entries $\mathcal{V}_{(d,1)}$, number of steps $K$, $G(t)$, Gumbel noise $\mathcal{G}$
3:Initialize: $x_{1}\leftarrow\mathcal{M}_{1}^{L}$, $t\leftarrow 1$.
4:for $k=1$ to $K$ do
5:     $t\leftarrow\frac{K-k+1}{K},s\leftarrow\frac{K-k}{K}$
6:     $\text{logits}\leftarrow f_{\theta}(x_{t},c)$
7:     $p_{\text{score}}\leftarrow\text{logits}+G(t)\cdot\mathcal{G}$
8:     $x_{\text{pred}}\leftarrow\text{argmax}(p_{\text{score}})$triangleright Predict-all
9:     $x_{s}\leftarrow\text{where}(x_{t}=[\text{m}],x_{\text{pred}},x_{t})$
10:     $p_{\text{conf}}\leftarrow p_{\text{score}}+G(t)\cdot\mathcal{G}$
11:     $m_{\text{re}}\leftarrow\text{argsort}(p_{\text{conf}})[1:L\cdot(1-\alpha_{s})]$
12:     $x_{s}\leftarrow\text{where}(m_{\text{re}},[\text{m}],x_{s})$triangleright Re-mask
13:end for
14:Return: fully unmasked sequence $x_{0}$
Algorithm 2 Rehash Sampling (Ours)
1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$.
2:Settings: codebook entries $\mathcal{V}_{(d,m)}$, number of steps $K$.
3:Initialize: $x_{1}\leftarrow\text{U}(\mathcal{M}_{m}^{L})$, $t\leftarrow 1$, $T^{1:K}$.
4:for $k=1$ to $K$ do
5:     $t\leftarrow T^{k},s\leftarrow T^{k+1}$
6:    # the rehash operation:
7:     $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$
8:     $\text{logits}\leftarrow f_{\theta}(x_{t},c)$
9:     $p_{0|t}\leftarrow\text{Softmax}(\text{logits})$
10:     $q_{s|t}\leftarrow\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\cdot p_{0|t}+\delta_{m[0]}\cdot\frac{1-\alpha_{s}}{1-\alpha_{t}}$
11:     $x_{\text{pred}}\leftarrow\text{Multnm.}(q_{s|t})$triangleright w/ masks
12:     $x_{s}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},x_{\text{pred}},x_{t})$
13:end for
14:Return: fully unmasked sequence $x_{0}$
In practice, we apply torch.multinomial (Multnm. in step 10) for low-discrepancy categorical sampling [18] with the merged probability of absorbing states  [^1] Instead of dividing $\alpha_{t|s}^{\rightarrow}$ and assigning these probabilities to each mask vocabulary. We merge the probabilities at step 9 in Alg. 2 to keep an overall noise sampling probability, as small values might be truncated.. The random nature of absorbing states inspires a rehash operation: we shuffle these tokens at the beginning of each step by $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$.
3.3 Discussion
Comparison with MVTM.
Masked visual token models (MVTMs) borrow the objective
$\mathcal{L}_{\text{MVTM}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t}),$ (19)
from masked language models [7] and predict on masked tokens with a maximum likelihood. Besides the reformulated corruption (Eq. 8) and reverse process (Eq. 17), ReDDiT differs in the following aspects: (i) the training objective (Eq. 4), which is derived from the principled DDM which provides better theoretical and empirical results with rehashing noise design. (ii) it can easily sample with arbitrarily discretized timeline, while MVTM couples training and inference, restricting its flexibility; (iii) the rehash sampler (Alg. 2) includes the absorbing states in categorical sampling with lower discrepancy, different from predict-remask sampler with time variant intensity $G(t)$ over Gumbel noise $\mathcal{G}$ (Alg. 1)  [^2] The logits corresponding to previously restored tokens’ indices are manually set to infinity for both methods, so that they will not be noised again in the following steps. This leads to an implementation of any-order auto-regressive model [26] if DDM’s decoded tokens per step is limited to $1$.. Gumbel-max suffers from numerical inaccuracy [49] and we detect that it worsens on large vocabulary (Fig. 3 with our reproduced results), which limits the potential of MVTM.
Relationship to DFM.
Discrete flow matching (DFM) [11] applies flow matching techniques to discrete structures and introduces a transition process based on masked tokens. Its training objective was initially designed as the masked token loss (19), and evolved to a time-weighted cross-entropy loss [34] for generalized diffusion paths, which is similar to ours. The similarity enables a direct comparison between the DFM sampler and our rehash sampler using the same trained model weights. We noticed that while the DFM sampler offers a refinement mechanism via token-wise updates, it generally requires more steps to reach optimal results. However, since the gradual decoding method is shared, we can integrate certain DFM steps into our sampling procedure for refinement. We leave this topic for an updated version of our essay.
Figure 3: Comparison of methods under limited samples. Left: Gumbel-max is theoretically equivalent to our method, yet it struggles to reflect the true distribution under limited sample passes. The multinomial approach captures the distribution more accurately. Right: our model achieves lower gFID across different sampling steps without tuning Gumbel-max, indicating more efficient and faithful sampling. (Best viewed in color)
4 Experiment
4.1 Implementation
Datasets.
[…]
Pre-processing.
Following the setting in LlamaGen [39], we apply the ten-crop augmentation on images, and use pre-trained tokenizers to convert them to discrete tokens. We pick IBQ-f16 [35] tokenizer as default for its scalable and promising performance in generation tasks, which uses a $16\times 16$ downsampling ratio and converts a $256\times 256$ image into 256 discrete tokens. The tokenizer has a codebook with 16384 entries. The LlamaGen-f16 (used in Tab. 3) and LlamaGen-f8 tokenizer [39] (used in Tab. 2) are also used for comparison with recent discrete generation methods. All tokenizers are used out-of-the-box without any tuning or modification.
Training with Representation Alignment.
Recent study [46] has shown that the alignment of intermediate representations between diffusion transformers and vision encoders accelerates training convergence of diffusion models. Accordingly, the alignment is designed as a regularization term with $\lambda=0.5$. We extract diffusion transformer’s n-th layer intermediate feature $\textbf{h}^{[n]}(x_{t})$ and align it with the original image’s dinov2-b [25] encoded features $f_{\text{enc}}(x_{0}^{\text{ori}})$. The intermediate features are projected by a small trainable MLP $h_{\varphi}$. The $\text{sim}(\cdot,\cdot)$ computes the mean of element-wise cosine similarity between embeddings, as
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DDM-linear}}+\lambda\mathcal{L}_{\text{RepA}},\ \ \mathcal{L}_{\text{RepA}}=-\mathbb{E}_{x,\ t}[\ \text{sim}(f_{\text{enc}}(x_{0}^{\text{ori}}),\ h_{\varphi}(\textbf{h}^{[n]}(x_{t})))\ ]\ .$ (20)
This alignment was proposed for continuous diffusion models, and we firstly validate that it’s also suitable for training discrete models in Tab. 1 (a).
Model Training.
The proposed model is based on DiT [28] architecture, with reference to its discrete prediction version [31]. 2D-RoPE [38] and min-SNR [47] are applied for training efficiency. The model is optimized using the AdamW optimizer with a cosine decay. Training is conducted for 500k iterations on 8 NVIDIA H100 GPUs with a global batch size 1024. Class-conditional training is enabled using class embeddings and a drop-rate of 0.1 for generation with classifier-free guidance.
4.2 Determining Noise Capacity
Figure 4: Comparison of different noise capacities. We re-implemented training with MVTM and ReDDiT design with the same training recipe (LlamaGen-f16 as visual tokenizer, with codebook size $16384$). The generation quality and representation alignment trends are visualized.
The reformulated discrete diffusion dynamics defines transitioning from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$. Under this setting, it is necessary to empirically determine the optimal value of $m$ for a fixed tokenizer with vocabulary size $d$, as the latent representations learned by VAEs are variant. We keep the training setup fixed and conduct experiments w.r.t. the noise capacity $m$. We also visualize $\mathcal{L}_{\text{RepA}}$, which captures the degree of representation alignment [46] within the transformer.
The alignment loss visualization shows that increasing the number of absorbing states introduces greater randomness, initially making predictions more difficult due to confusion with valid tokens. However, this gap narrows as training progresses, and the model converges to a similar alignment lower bound, suggesting effective representation learning across different configurations.
As shown in Fig. 4 (left), generation quality improves with increasing noise capacity initially. The LlamaGen-f16 tokenizer achieves peak performance at $m=128$, while the IBQ tokenizer performs best at $m=1024$. We attribute this to the codebook design: the lower dimensional LlamaGen-f16 codebook produces more compact latents, which also determines its smaller noise endurance.
4.3 Ablation Study
Table 1: Ablated Design choices. ReDDiT-L is trained for 100k iters. Final setting denoted in gray.
(a) (a) RepA choices
Layer RepA gFID
- $\times$ 7.24
4 5.72
8 5.41
12 5.82
16 6.12
(b) (b) General model design
Configuration gFID Prec. Rec.
MVTM (19) + RepA 6.83 0.75 0.39
    objective (20) 6.25 0.76 0.43
+ 2D-RoPE + min-SNR 5.97 0.78 0.44
+ rehash sampler 5.53 0.81 0.51
+ DFM refine 5.40 0.81 0.52
(c) (c) Sampling timeline
Steps Timeline gFID
20 linear 7.18
32 linear 6.43
20 arccos 5.04
20 square 7.39
20 cosine 4.91
Unless specified, all the models are trained on ImageNet $256\times 256$ under the default settings for 100k iterations for fair comparison. We use a constant guidance scale of 2.0 and 20 steps for generation, and report gFID $\downarrow$ and IS $\uparrow$ computed on 50K samples. Precision (Prec.$\uparrow$) and Recall (Rec.$\uparrow$) are also reported in general design for direct diversity comparison.
General Design.
We ablate the general choices of ReDDiT, which starts with a re-trained MVTM baseline methods (with LlamaGen-f16 and RepA for faster convergence as default) in Tab. 1. We also tested RepA supervision on different layers of ReDDiT. The results are consistent with the original study [46], which indicates that adding RepA objective at the 8-th of 24 layers results in best convergence for L-model. The objective and modifications are conducted with re-training. We replace the MVTM sampler with the proposed rehash sampler and DFM refinement for further improvement.
Sampling Timeline.
Figure 5: Illustration of discretized timeline with $K=7$. The slow-to-fast sampling works better than linear schedules.
Recovering complete information from noise remains critical to diffusion-based models [22, 42]. Recent work shows MVTM’s non-linear scheduler for training is less critical when using high-capacity tokenizers. Evidence of time-invariance in DDMs [31, 36] further supports decoupling training from sampling. In our experiments, a linear scheduler—with constant signal-to-noise ratio decay—yields optimal training dynamics. Among the timeline discretization tested, Fig. 5, the cosine schedule is employed for our ReDDiT model for its optimal performance in Tab. 1.
4.4 Performance Comparison
We compare the proposed ReDDiT model with other generative models on the ImageNet-1K $256\times 256$ in Tab. 2. The IBQ tokenizer is used for the default L and XL models. We also utilize LlamaGen-f8 with 128 noise capacity to evaluate its high-resolution potentials (noted as ReDDiT-XLf8). We use a linear increasing guidance following the common practice of  [10].
Table 2: Performance comparison on class-conditional ImageNet 256$\times$256. Look-up free quantizers are beyond the scope of this paper. ft. indicates that the decoder is fine-tuned for quantized latents. Wall-clock inference time relative to ReDDiT-XL is reported.
Type Model Tokenizer Generator
[…]
Diff. LDM-4 [29] 4096×3 - 3.60 247.7 400M 250
DiT-XL/2 [28] 1024×4 - 2.27 278.2 675M 250 18
MDTv2 [10] 1024×4 - 1.58 314.7 676M 250 >18
MAR-H [20] 256×16 - 1.55 303.7 943M 256 10
MVTM MaskGIT [4] 256 1024 6.18 182.1 227M 8 0.2
TiTok-S-128ft. [45] 128 4096 1.97 281.8 287M 64 1.6
AR Taming-VQGAN [9] 256 1024 15.78 74.3 1.4B 256 8
RQ-Transformer [17] 256 16384 7.55 134.0 3.8B 64 8.5
ViT-VQGAN [43] 1024 8192 4.17 175.1 1.7B 1024 >10
LlamaGen-3B [39] 576 16384 2.18 263.3 3.1B 576 20
RandAR-XXL [27] 512 16384 2.15 322.0 1.4B 88 4
VAR-d30 [41] 680 4096 1.97 334.7 2.0B 10 0.5
DDM ITM [16] 1024 16384 5.30 183.0 546M 100 3
ReDDiT-L (ours) 256 16384 2.13 294.7 346M 20 0.5
ReDDiT-XL (ours) 256 16384 1.74 313.6 675M 32 1
ReDDiT-XLf8 (ours) 1024 16384 1.61 318.5 675M 64 2
Table 3: Comparison of models with the same tokenizer. Reconstruction FID (rFID) indicates the tokenizer’s reconstruction quality from its quantized codes. Dim denotes codebook dimension. AR model’s gFID are indexed from their original report.
Model VQ Tokenizer Info. Generator
[…]
LlamaGen-LAR [39] LlamaGen-f16 [39] 2.19 8 343M 3.80
RandAR-LAR [27] 343M 2.55
OursDDM(ReDDiT-L) 346M 2.41
IBQ-BAR [35] IBQ-tokenizer [35] 1.37 256 343M 2.88
OursDDM(ReDDiT-L) 346M 2.13
Generation Quality.
As shown in Tab. 2, ReDDiT achieves the best performance among the compared discrete models. It outperforms the baseline (MaskGIT [4]) with significant margins (gFID: 2.13 vs 6.18 and IS: 294.7 vs. 182.1). It also outperforms the recent DDM method [16] and TiTok-S-128 [45], which is extensively fine-tuned on quantized latents. Compared with continuous diffusion models, ReDDiT exhibits higher efficiency and on-par performance, showing great potential for discrete generation. Note that the performance is achieved with a codebook size of 16384, validating ReDDiT’s effectiveness for large-vocabulary codebooks.
Efficiency.
ReDDiT is born with the high-efficiency advantage of discrete diffusion models. As shown in Tab. 2, the inference time of ReDDiT is slightly longer than MaskGIT, while the performance is overwhelming. With only 20 steps it achieves a competitive performance which diffusion models uses more than 250 steps to achieve. It also demonstrates comparable results with significantly smaller model size than recent VAR [41] method.
Besides the major comparison, we also conduct an experiment that utilizes the identical tokenizer in previous auto-regressive (AR) works and validate our method’s effectiveness. As can be seen in Tab. 3, ReDDiT outperforms AR methods in generation tasks across different tokenizers.
4.5 Qualitative Results
Figure 6 presents representative class-conditional samples generated by the proposed ReDDiT model. The outputs across diverse image classes consistently exhibit both high fidelity and diversity. We further demonstrate ReDDiT’s editing capability in Figure 7, highlighting its bi-directional perceptual competence. Following MaskGIT [4], we replace a region of the input image with noise tokens and employ the same generation pipeline to inpaint the missing content, conditioned on a class label $c$.
Figure 6: Upper: Comparison between MVTM and our method on generated images. Below: Class-conditional generation samples of ReDDiT on ImageNet $256\times 256$.
Figure 7: Class-conditional generation and in-painting samples of ReDDiT on ImageNet $256\times 256$.
5 Conclusion
We proposed ReDDiT, a discrete visual generative model built upon a principled discrete diffusion architecture with novel noise designs and efficient sampling strategies. Our key contribution lies in the integration of rehashing noise and samplers, which together ensure both diversity and low discrepancy throughout the generative process. By introducing rehashing noise, ReDDiT enriches the potential paths that latent variables can traverse during training, enhancing representational capacity. Extensive experiments demonstrate that discrete generative models can achieve performance on par with their continuous counterparts while offering significantly improved efficiency. This work paves a promising way for discrete generative modeling and offers fresh insights toward unifying visual and language generation—a path we leave for future exploration.

### v3

@@
 Abstract
-Discrete diffusion models are gaining traction in the visual generative area for their efficiency and compatibility.
-However, the pioneered attempts still fall behind the continuous counterparts, which we attribute to the noise (absorbing state) design and sampling heuristics.
-In this study, we propose the rehashing noise framework for discrete diffusion transformer (termed ReDDiT), to extend absorbing states and improve expressive capacity of discrete diffusion models.
+In the visual generative area, discrete diffusion models are gaining traction for their efficiency and compatibility.
+However, pioneered attempts still fall behind their continuous counterparts, which we attribute to noise (absorbing state) design and sampling heuristics.
+In this study, we propose a rehashing noise approach for discrete diffusion transformer (termed ReDDiT), with the aim to extend absorbing states and improve expressive capacity of discrete diffusion models.
 ReDDiT enriches the potential paths that latent variables traverse during training with randomized multi-index corruption.
-The derived rehash sampler, which reverses the randomized absorbing paths, guarantees the diversity and low discrepancy of the generation process.
+The derived rehash sampler, which reverses the randomized absorbing paths, guarantees high diversity and low discrepancy of the generation process.
 These reformulations lead to more consistent and competitive generation quality, mitigating the need for heavily tuned randomness.
-Experiments show that ReDDiT significantly outperforms the baseline (reducing gFID from 6.18 to 1.61) and is on par with the continuous counterparts with higher efficiency.
+Experiments show that ReDDiT significantly outperforms the baseline model (reducing gFID from 6.18 to 1.61) and is on par with the continuous counterparts.
+The code and models will be publicly available.
 1 University of Chinese Academy of Sciences 2 China Mobile Research Institute
 matianren18@mails.ucas.ac.cn qxye@ucas.ac.cn
-Project Page: https://github.com/martian422/ReDDiT
 1 Introduction
-Diffusion has been a highly competitive approach for generative workloads [8, 30, 21], offering strong bidirectional perception and well-structured mechanisms [48] for global control over content.
-In the continuous domain, diffusion transformers (DiTs) [28], which progressively refine image latents from Gaussian noise, have achieved impressive and scalable results.
-Despite the popularity of DiTs, the community shows a growing interest in discrete diffusion models in recent years [16, 40].
-This trend is based on their practical advantages, e.g., compatibility with language models for the indexable codebooks and efficiency for the ability to generate multiple tokens at each inference.
-Early endeavors [4, 5, 12] mainly pursue efficiency through integrating visual tokenizers and BERT-style [mask] tokens [7].
-Recent studies [2, 45] improved the generation quality, demonstrating great potential of visual generation in discrete domain.
-Despite the progress, discrete diffusion methods remain lagging behind their continuous counterparts.
-State-of-the-art approaches, e.g., masked visual token models (MVTMs) [45], is puzzeld by its mask design and predict-remask sampler, which restricts the model’s expressive capacity and makes the prediction performance sensitive to adaptions even with extensive training, Fig.
+Diffusion has been a competitive approach for generative workloads Dhariwal & Nichol (2021); Rombach et al.
+(2022b); Li et al.
+(2024), offering strong bidirectional perception and well-structured mechanisms Zhang et al.
+(2023) for global control over content.
+Within the continuous domain, diffusion transformers (DiTs) Peebles & Xie (2023), which progressively refine image latents from Gaussian noise, have achieved impressive and scalable results.
+Recently, the community shows a growing interest in discrete diffusion models Hu & Ommer (2024); Swerdlow et al.
+(2025), which is based on their practical advantages, e.g., compatibility with language models for the indexable codebook, and efficiency for predicting multiple tokens at each inference.
+Early endeavors Chang et al.
+(2022; 2023); Gu et al.
+(2022) pursue efficiency through integrating visual tokenizers and BERT-style [mask] tokens Devlin et al.
+(2019).
+Recent studies Bai et al.
+(2025); Yang et al.
+(2025) improved the generation quality, demonstrating great potential of discrete diffusion.
+Despite the progress, the performance of discrete diffusion methods remains lagging behind their continuous counterparts.
+Representative approaches, e.g., masked visual token models (MVTMs) Chang et al.
+(2022); Yu et al.
+(2023), are puzzled by the mask design and confidence-based re-mask sampler, which restricts model’s expressive capacity and makes prediction sensitive to adaptions given extensive training, Fig.
 1(upper).
-Moreover, when paired with large-vocabulary codebooks from high-fidelity modern tokenizers, these models encounter challenges such as slower sampling speeds and numerical inaccuracy [49].
-Figure 1: Comparison of ReDDiT with the baseline discrete model (MVTM).
+Moreover, when paired with large-vocabulary codebooks from high-fidelity modern tokenizers, they encounter challenges such as slower sampling speeds and numerical inaccuracy Zheng et al.
+(2024).
+Figure 1: Comparison the baseline discrete model (MVTM) with ReDDiT.
 MVTMs rely on score-based remasking strategies with Gumbel-max to sample from logits, which leads to lower token diversity and suboptimal token selection.
-In contrast, ReDDiT introduces a principled, low-discrepancy rehashing mechanism that leverages softmax-based probabilities, enabling diverse, high-quality sampling through a learned distribution.
-(Best viewed in color)
-To address these limitations, we propose two key hypotheses.
-First, while previous methods learn to recover plausible tokens from a monotonous [mask] canvas, this noise design may not be well-suited for discrete visual generation.
-In continuous diffusion, Gaussian noise is used to progressively degrade the input to learn a smooth distribution shift  [15, 22].
-Discrete masking mimics this paradigm by collapsing all masked tokens into a single absorbing state, which however, lacks the variability of Gaussian noise, in terms of both vocabulary richness and latent diversity.
-Consequently, the discrete process offers a far coarser signal, which limits its ability to capture and represent diverse data distributions [33, 1].
-Moreover, while continuous diffusion models introduce stochasticity at every inference step through noise injection, discrete unmasking is inherently binary: tokens are either masked or deterministically decoded.
-This rigid mechanism further constrains the flexibility of sample refinement during generation.
-Second, the predict-remask sampler of MVTMs introduces a form of handcrafted randomness, which is often implemented using Gumbel-max, to approximate sampling diversity.
-However, this sampler may compromise the probabilistic fidelity of the generation, and the need to carefully balance the token numbers decoded per step to mitigate accumulation errors leads to redundant sampling passes.
-As a result, the Gumbel-max has evolved to a heavily tuned time variant trick, leading to unstable performance, particularly when scaled to large-vocabulary codebooks.
-The above factors, rather than quantization alone, attribute to the performance gap between discrete and continous models.
+In contrast, ReDDiT introduces a systematic, low-discrepancy rehashing mechanism that leverages softmax-based probabilities, enabling diverse, high-quality sampling through a learned distribution.
+(This figure is best viewed in color)
+To address these limitations, we first propose two hypotheses.
+First, while discrete methods learn to recover plausible tokens from a monotonous [mask] canvas, the used noise design may not be well-suited for discrete visual generation.
+In continuous diffusion, Gaussian noise is used to progressively degrade the input to learn a smooth distribution shift  Ho et al.
+(2020); Lu et al.
+(2022).
+Discrete masking mimics this paradigm by collapsing all masked tokens to a single absorbing state, which, however, lacks the variability of Gaussian noise, in terms of both vocabulary richness and latent diversity.
+Consequently, the discrete process offers a far coarser signal, which limits its ability to represent diverse data distributions Santos et al.
+(2023); Austin et al.
+(2021).
+Moreover, while continuous diffusion models introduce stochasticity at every inference step through noise injection, discrete unmasking is inherently binary: tokens are either masked or deterministically decoded, Fig.
+1(upper).
+This rigid mechanism constrains the flexibility of sample refinement during generation.
+Second, the confidence-based re-mask sampler of MVTMs introduces a form of handcrafted randomness, which is implemented through Gumbel-max, to approximate sampling diversity.
+Unfortunately, this sampler compromises the probabilistic fidelity of generation, and the need to carefully balance token numbers decoded per step (for mitigating accumulation errors) leads to redundant sampling passes.
+As a result, Gumbel-max has evolved to a heavily tuned time variant trick with unstable performance, particularly when scaled to large-vocabulary codebooks.
+The above factors, rather than quantization alone, induce the performance gap between discrete and continuous models.
 In this study, we propose a discrete diffusion model with an elaborate rehashing noise design, Fig.
@@
 Specifically, we expand the masks to multiple indices along with the codebook and randomize them during data corruption.
-A rehash sampler is also derived with principled discrete diffusion theories to reverse the diffusion path for generation, guaranteeing the diversity and low discrepancy of the sampling process.
-We demonstrate that this rehashed noise helps learning a superior and flatter lower error bound while eliminating reliance to hyper-parameterized randomness during sampling.
-We further revisit the commonly used discrete diffusion objective (e.g., the masked cross-entropy loss) and update it with empirical modifications.
-By adopting an improved ELBO [31, 36] with representation alignment (RepA) [46] loss, we achieve better training efficiency and generation quality.
-Moreover, ReDDiT aligns with recent advances in discrete flow matching [11, 34], enabling token refreshment during sampling without training post-correction models [19].
-2 Related Work
-Diffusion Models.
-Diffusion models [15, 37] have emerged as a powerful class of generative methods that learn data distributions by reversing a gradual noising process over time.
-These models are primarily designed for continuous domains such as images [8, 10, 28], defining a forward process that transforms data $x_{0}$ into noise $x_{1}$: $x_{t}\sim\mathcal{N}(\sqrt{\alpha_{t}}x_{0};(1-\alpha_{t})\mathbf{I})$ where $\alpha_{t}$ controls the noise schedule.
-The generative (reverse) process learns a denoising model $p_{\theta}(x_{s}\mid x_{t})$, often parameterized via a neural network $\theta$ to predict either noise or clean data.
-Discrete Diffusion Models.
-Discrete diffusion has been previously governed by masked visual token models (MVTMs) [4, 5, 12, 44, 45].
-This model leverages a BERT-style [mask] token to corrupt the tokenized image sequence and trained the network with a simple cross-entropy loss on masked tokens, resulting in a score-based prediction.
-It generates tokens in a non-autoregressive fashion, by remasking the tokens with least scores at each inference as depicted in Alg.
-1.
-Recent studies unlocked the principled discrete diffusion model (DDM) [31, 36] and discrete flow-matching (DFM) [11, 34], which adapt the Markov chain theory, enabling generation over text [26, 23], molecules[34], and other discrete representations [1, 24].
-Unlike MVTMs, the principled DDM and DFM mostly derive a time-weighted cross-entropy loss to supervise the training procedure and apply a gradual unmasking method based on probabilities.
+A rehash sampler is also derived with principled discrete diffusion theories to reverse the diffusion path for generation, guaranteeing high diversity and low discrepancy of the sampling process.
+We demonstrate that this rehashed noise facilitates learning a superior and regularized expressiveness, while eliminating reliance to hyper-parameterized randomness during sampling.
+We further revisit the commonly used discrete diffusion objective and update it with empirical modifications.
+By adopting an improved ELBO Sahoo et al.
+(2024); Shi et al.
+(2024) with representation alignment (RepA) Yu et al.
+(2025) loss, we optimize the training efficiency and substantially improve the generation quality of discrete generative models.
+Moreover, ReDDiT aligns with recent advances in discrete flow matching Gat et al.
+(2024); Shaul et al.
+(2024), enabling token refreshment during sampling without training post-correction models Lezama et al.
+(2022).
 2 Methodology
 For self-containment, we first review the DDM theory in Sec.
-3.1.
-We reformulate its diffusion dynamics and introduce rehashing noise for ReDDiT in Sec.
-3.2.
-Connection and comparison of ours and other discrete diffusion models are discussed in Sec.
-3.3.
+2.1.
+We then reformulate its diffusion dynamics and introduce rehashing noise for ReDDiT in Sec.
+2.2.
+We finally discuss connection and comparison with other discrete diffusion models in Sec.
+2.3.
 2.1 Preliminary: Discrete Diffusion Model
-DDM defines a forward process over discrete variables by gradually corrupting the image tokens into absorbing states (masks) through a continuous-time Markov process.
-Assume that the data consists of tokens from a finite vocabulary $\mathcal{V}$, $x\in\mathcal{V}^{L}$ represents a sequence of tokens ($e.g.$, an image tokenized into indices) with length $L$.
+DDM defines a forward process over discrete variables by gradually corrupting the image tokens to absorbing states (masks) through a continuous-time Markov process.
+Assume that the data consists of tokens from a finite vocabulary $\mathcal{V}$.
+$x\in\mathcal{V}^{L}$ is a sequence of tokens ($e.g.$, an image tokenized into indices) with length $L$.
 We denote the clean data as $x_{t=0}$ ($x_{0}$ for short), and noise it gradually as $t\rightarrow 1$.
@@
 Let $\alpha_{t}$ be the noise scheduler (a monotonically decreasing survival function that satisfies $\alpha_{0}=1,\alpha_{1}=0$ ).
-For $0\leq s<t\leq 1$, denote $\alpha_{t|s}^{\leftarrow}=\alpha_{t}/\alpha_{s}$, the forward corruption process is governed by a continuous-time transition kernel $q(x_{t}^{i}\mid x_{s}^{i})$ on the $i$-th element, as
-$q(x_{t}^{i}\mid x_{s}^{i})=\begin{cases}1-\alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=\mathbf{m},x_{s}^{i}\neq\mathbf{m},\\ \alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}\neq\mathbf{m},\\ 1,&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}=\mathbf{m},\\ 0,&\text{otherwise}.\end{cases}$ (1)
-Given the transition kernel, the corrupted data distribution at time $t$ is determined  [31] as
-$x_{t}\sim\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{m}^{L}$ (2)
-The generative model learns the reverse process $p_{\theta}(x_{s}\mid x_{t})$, which denoises sample $x_{t}$ at arbitrary time $t\in(0,1]$ to a less noised state $x_{s}$ at time $s<t$.
-Denote $\delta(x_{t}^{i},\textbf{m})$ as the indicator function that only computes on masked tokens, and $\alpha_{t}^{\prime}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$, the learning objective is derived  [36, 31] as
-$\mathcal{L}_{\text{DDM}}=-\mathbb{E}_{x_{0},\ x_{t}}\int_{t=0}^{t=1}[\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t})]\mathrm{d}t\ .$ (3)
-For a linear scheduler, Eq. 3 is simplified via variable substitution [36, 23] to an equivalent form:
-$\mathcal{L}_{\text{DDM-linear}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}[\frac{1}{t}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t})]\ .$ (4)
+For $0\leq s<t\leq 1$, the forward corruption process is governed by a continuous-time transition kernel $q(x_{t}^{i}|x_{s}^{i})$ at the $i$-th element, as
+$q(x_{t}^{i}|x_{s}^{i})=\begin{cases}1-\alpha_{t|s},&\text{if }x_{t}^{i}=\mathbf{m},x_{s}^{i}\neq\mathbf{m}\\ \alpha_{t|s},&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}\neq\mathbf{m}\\ 1,&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}=\mathbf{m}\\ 0,&\text{otherwise}\end{cases},\quad\alpha_{t|s}=\frac{\alpha_{t}}{\alpha_{s}}.$ (1)
+Denoting $q$ as the transition kernel and $\text{Cat}(\cdot;\pi)$ the categorical distribution determined by probability $\pi$, the corrupted data distribution at time $t$ is written as
+$x_{t}\sim q(x_{t}|x_{0}),q(x_{t}|x_{0})=\text{Cat}(x_{t};\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{m}^{L}).$ (2)
+The generative model learns the reverse process $p_{\theta}(x_{s}|x_{t})$, which denoises sample $x_{t}$ at arbitrary time $t\in(0,1]$ to a less noised state $x_{s}$ at time $s<t$.
+Denoting $\delta(x_{t}^{i},\textbf{m})$ as the indicator function that only computes on masked tokens, and $\alpha_{t}^{\prime}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$, the learning objective is derived  Shi et al.
+(2024) as
+$\mathcal{L}_{\text{DDM}}=-\mathbb{E}_{x_{0},\ x_{t}}\int_{t=0}^{t=1}[\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}|x_{t})]\mathrm{d}t\ .$ (3)
+For a linear scheduler, Eq. 3 is simplified via variable substitution Sahoo et al. (2024) to an equivalent form, as
+$\mathcal{L}_{\text{DDM-linear}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}[\frac{1}{t}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}|x_{t})]\ .$ (4)
 For conditional generation, class information $c$ (e.g., labels or text prompts) is introduced to the denoising model as additional input.
-Following classifier-free guidance [14], the model is trained with a random drop of labels, and the prediction is interpolated at inference:
+Following classifier-free guidance Ho & Salimans (2022), the model is trained with a random drop of labels, and the prediction is interpolated at inference, as
 $\hat{p}_{\theta}(x_{t},c)=\ p_{\theta}(x_{t},\varnothing)+w\cdot(\,p_{\theta}(x_{t},c)-\ p_{\theta}(x_{t},\varnothing)),$ (5)
-where $w\geq 0$ controls the guidance strength.
+where $\varnothing$ is the dropped label and $w\geq 0$ controls the guidance strength.
+2.2 Discrete Diffusion with Rehashing Noise
+The ordinal structure inherent in discrete data provides a valuable inductive bias for designing transition kernels in diffusion dynamics.
+Prior studies Austin et al.
+(2021); Campbell et al.
+(2022) show that assigning higher transition probabilities to neighboring pixel values—forming a discrete Gaussian-like noise—outperforms the single absorbing state approach on pixel-level datasets like CIFAR-10.
+However, when using visual tokenizers, the structure of discretized latents is learned rather than pre-defined, making such ordinal assumptions inapplicable.
+This insight motivates us to extend conventional mask tokens to a set of indices, and reverse the diffusion path with noise rehashing.
+This design allows the model to optimize its embedding space during training, enhancing its ability to model flexible and data-driven noise structures.
+We visualize the learned distributions in Fig.
+2 (right).
 Figure 2: Visualization of pixel and latent spaces.
@@
 3D t-SNE plot (right) is used solely to separate valid and noise tokens for clustering illustration.
-3.2 Discrete Diffusion with Rehashing Noise
-The ordinal structure inherent in discrete data provides a valuable inductive bias for designing transition kernels in diffusion dynamics.
-Prior studies [1, 3] show that assigning higher transition probabilities to neighboring pixel values—forming a discrete Gaussian-like noise—outperforms the single absorbing state approach on pixel-level datasets like CIFAR-10.
-However, when using visual tokenizers, the structure of discretized latents is learned rather than pre-defined, making such ordinal assumptions inapplicable.
-This insight motivates us to extend the conventional mask token into a set of indices, and reverse the diffusion path with noise rehashing.
-Such a procedure allows the model to optimize its embedding space during training, enhancing its ability to model flexible and data-driven noise structures.
-We visualize the learned distributions in Fig.
-2 (right).
 Reformulation.
 Given $d$ categories, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be its one-hot vector where the $i$-th value is $1$.
-We denote =$\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution (known as $d$-simplex), and a basis for absorbing states with capacity $m$: =$\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$.
+We denote $\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution (known as $d$-simplex), and a basis for absorbing states with capacity $m$: $\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$.
 The sum of $\mathcal{E}$ and $\mathcal{M}$ can be denoted as
@@
 $\mathcal{E}_{d}=\left\{\mathbf{v}_{(i,0)}\in\mathcal{V}_{(d,m)}\,\middle|\,i=1,\ldots,d\right\},\ \mathcal{M}_{m}=\left\{\mathbf{v}_{(0,j)}\in\mathcal{V}_{(d,m)}\,\middle|\,j=1,\ldots,m\right\}.$ (7)
-To exploit visits across all the possible paths, we rewrite the transition kernel defined by Eq. 1 into
+To exploit visits across all the possible paths, we rewrite the transition kernel defined by Eq. 1 as
+$q(x_{t}^{i}|x_{s}^{i})=\begin{cases}1-\alpha_{t|s},&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\notin\mathcal{M}_{m}\\ \alpha_{t|s},&\text{if }x_{t}^{i}=x_{s}^{i},\ x_{s}^{i}\notin\mathcal{M}_{m}\\ 1/m,&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\in\mathcal{M}_{m}\\ 0,&\text{otherwise}.\end{cases}$ (8)
+With above definitions, we reformulate the diffusion process of $x$ as a transition from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$.
+We train the model by feeding it with corrupted data, of which the distribution is inferred as $x_{t}\sim\text{Cat}(x_{t};\alpha_{t}x_{0}+(1-\alpha_{t})\text{U}(\mathcal{M}_{m}^{L}))$, where $\text{U}(\mathcal{M}_{m}^{L})$ is the uniform distribution upon $\mathcal{M}_{m}^{L}$.
+Rehash Sampling.
+To generate a sequence of length $L$, the reverse process starts with $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$.
+The subsequent latents $x_{t}$ are generated by discretizing the reverse timeline $T$ to $K$ steps.
+We denote this schedule as $T^{1:K+1}$ such that $T^{1}=1$ and $T^{K+1}=\varepsilon$, with $\varepsilon$ being an arbitrarily small positive constant.
+The reverse process is deduced from the formulation, as
+[…]
+Comparing with MVTM sampler in Alg.
+1, our rehash sampler is shown in Alg.
+2.
+Similar to MDLM Sahoo et al.
+(2024), we apply torch.multinomial (Multnm.
+in step 10) for low-discrepancy [^1] Instead of dividing $\alpha_{s|t}$ and assigning these probabilities to each mask vocabulary.
+We merge the probabilities at step 9 to keep an overall noise sampling probability, as small values might be truncated.
+categorical sampling.
+Algorithm 1 MVTM Sampling
+1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$,
+2:Settings: number of steps $K$, $G(t)$, $\mathcal{G}$
+3:Initialize: $x_{1}\leftarrow\mathcal{M}_{1}^{L}$, $t\leftarrow 1$.
+4:for $k=1$ to $K$ do
+5:  $t\leftarrow\frac{K-k+1}{K},s\leftarrow\frac{K-k}{K}$
+6:  $p_{\text{score}}\leftarrow f_{\theta}(x_{t},c)+G(t)\cdot\mathcal{G}$
+7:  $x_{\text{pred}}\leftarrow\text{argmax}(p_{\text{score}})$ $\triangleright$ Predict-all
+8:  $x_{s}\leftarrow\text{where}(x_{t}=[\text{m}],x_{\text{pred}},x_{t})$
+9:  $p_{\text{conf}}\leftarrow p_{\text{score}}+G(t)\cdot\mathcal{G}$
+10:  $m_{\text{re}}\leftarrow\text{argsort}(p_{\text{conf}})[1:L\cdot(1-\alpha_{s})]$
+11:  $x_{s}\leftarrow\text{where}(m_{\text{re}},[\text{m}],x_{s})$ $\triangleright$ Re-mask
+12:end for
+13:Return: fully unmasked sequence $x_{0}$
+Algorithm 2 Rehash Sampling
+1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$.
+2:Settings: number of steps $K$.
+3:Initialize: $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$, $t\leftarrow 1$, $T^{1:K}$.
+4:for $k=1$ to $K$ do
+5:  $t\leftarrow T^{k},s\leftarrow T^{k+1}$
+6:  $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$
+7:  $p\leftarrow\pi_{\theta}(x_{t},c)$
+8:  $q_{s|t}\leftarrow\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\cdot p+\delta\cdot\frac{1-\alpha_{s}}{1-\alpha_{t}}$
+9:  $x_{\text{pred}}\leftarrow\text{Multnm.}(q_{s|t})$ $\triangleright$ w/ masks
+10:  $x_{s}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},x_{\text{pred}},x_{t})$
+11:end for
+12:Return: fully unmasked sequence $x_{0}$
+The random nature of absorbing states inspires a rehash operation: we shuffle these tokens at the beginning of each step by $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$.
+Proof to Eq.9 is included in Appendix.
+C.
+2.3 Discussion
+Figure 3: Sampler comparison.
+Left: Gumbel-max is theoretically equivalent to our method, yet it struggles to reflect the true distribution under limited sample passes.
+The multinomial approach captures the distribution more accurately.
+Right: our model achieves lower gFID across different sampling steps without tuning Gumbel-max, indicating more efficient and faithful sampling.
+a, b, c refer to three uniformly sampled $G(t)$ set for MVTM sampling.
+See supplementary for experimental codes.
+(This figure is best viewed in color)
+Comparison with MVTM.
+Masked visual token models (MVTMs) borrow the objective
+$\mathcal{L}_{\text{MVTM}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t}),$ (10)
+from masked language models Devlin et al.
+(2019) and predict on masked tokens with a maximum likelihood.
+Besides the reformulated corruption (Eq.
+8) and reverse process (Eq.
+9), ReDDiT differs in the following aspects: (i) the training objective (Eq.
+4), which is derived from DDM, providing better theoretical and empirical results.
+(ii) it can easily sample with a arbitrarily discretized timeline, while MVTM couples training and inference, restricting its sampling flexibility; (iii) the rehash sampler (Alg.
+2) includes absorbing states in categorical sampling with lower discrepancy, different from MVTM’s predict-remask sampler with time variant intensity $G(t)$ over Gumbel noise $\mathcal{G}$ (Alg.
+1)  [^2] The logits corresponding to previously restored tokens’ indices are manually set to infinity for both methods, so that they will not be noised again in the following steps.
+This leads to an implementation of any-order auto-regressive model Ou et al.
+(2024) if DDM’s decoded tokens per step is limited to $1$..
+Gumbel-max suffers from numerical inaccuracy Zheng et al.
+(2024) and we noitice that it becomes worse on large vocabulary (Fig.
+13 with our reproduced results), which limits MVTM’s potential.
+Relationship to DFM.
+Discrete flow matching (DFM) Gat et al.
+(2024) introduces a transition process based on masked tokens.
+Its training objective was initially designed as the masked token loss ( 10), and evolved to a time-weighted cross-entropy loss Shaul et al.
+(2024) for generalized diffusion paths, which is similar to ours.
+The similarity enables a direct comparison between the DFM sampler and our rehash sampler using the same trained model weights.
+We notice that it generally requires more steps to reach optimal results, as the DFM sampler offers a refinement mechanism via token-wise updates.
+Since the gradual decoding method is shared, we can integrate certain DFM steps into our sampling procedure for refinement.
+This leads to $\sim$ 0.1 gFID improvement on ImageNet-1K.
+Refer to Appendix D for details.
+3 Experiment
+3.1 Implementation
+Datasets.
+[…]
+Pre-processing.
+Following the setting in LlamaGen Sun et al.
+(2024), we apply the ten-crop augmentation on images, and use pre-trained tokenizers to convert them to discrete tokens.
+We pick IBQ-f16 Shi et al.
+(2025) tokenizer as default for its scalable and promising performance in generation tasks, which uses a $16\times 16$ downsampling ratio and converts a $256\times 256$ image into 256 discrete tokens.
+The tokenizer has a codebook with 16384 entries.
+The LlamaGen-f16 (used in Tab.
+2) and LlamaGen-f8 tokenizer Sun et al.
+(2024) (used in Tab.
+1) are also used for comparison with recent discrete generation methods.
+All tokenizers are used out-of-the-box without modification.
+Representation Alignment.
+Recent study Yu et al.
+(2025) has shown that the alignment of intermediate representations between diffusion transformers and vision encoders accelerates training convergence of diffusion models.
+Accordingly, the alignment is designed as a regularization term with $\lambda=0.5$.
+We extract diffusion transformer’s 8-th layer intermediate feature $\textbf{h}^{[n]}(x_{t})$ and align it with the original image’s dinov2-b Oquab et al.
+(2023) encoded features $f_{\text{enc}}(x_{0}^{\text{ori}})$.
+The intermediate features are projected by a small trainable MLP $h_{\varphi}$.
+The $\text{sim}(\cdot,\cdot)$ computes the mean of element-wise cosine similarity between embeddings, as
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DDM-linear}}+\lambda\mathcal{L}_{\text{RepA}},\ \ \mathcal{L}_{\text{RepA}}=-\mathbb{E}_{x,\ t}[\ \text{sim}(f_{\text{enc}}(x_{0}^{\text{ori}}),\ h_{\varphi}(\textbf{h}^{[n]}(x_{t})))\ ]\ .$ (11)
+This alignment was proposed for continuous diffusion models, and we firstly validate that it’s also suitable for training discrete models.
+However, from our observation, as a training acceleration technique, RepA does not provide relative performance gain if training sufficiently (like for 1M steps as most diffusion models do) for discrete latents.
+We only use RepA to improve training efficiency and probe the inner dynamics through training as in Fig.
+4.
+Training and Evaluation.
+The proposed model is based on DiT Peebles & Xie (2023) architecture, with reference to its discrete prediction version Sahoo et al.
+(2024).
+2D-RoPE Su et al.
+(2024) and min-SNR Zhang & Sennrich (2019) are applied for training efficiency.
+The model is optimized using the AdamW optimizer with a cosine decay.
+Training is conducted for 500k iterations on 8 NVIDIA H100 GPUs with a global batch size 1024.
+Class-conditional training is enabled using class embeddings and a drop-rate of 0.1 for generation with classifier-free guidance.
+Details are provided in Appendix E.
+3.2 Performance and Comparison
+We compare the proposed ReDDiT model with other generative models on the ImageNet-1K $256\times 256$ in Tab.
+1.
+The IBQ tokenizer is used for the default L and XL models.
+We also utilize LlamaGen-f8 with 128 noise capacity to evaluate its high-resolution potentials (noted as ReDDiT-XLf8).
+We use a linear increasing guidance following the common practice of  Gao et al.
+(2023).
+Table 1: Performance comparison on class-conditional ImageNet $\times256$.
+Look-up free quantizers are beyond the scope of this paper.
+ft.(in gray) indicates that the decoder is fine-tuned for quantized latents.
+Wall-clock inference time relative to ReDDiT-XL is reported.
+Type Model Tokenizer Generator
+[…]
+Diff. LDM-4 Rombach et al. (2022a) 4096×3 - 3.60 247.7 400M 250
+DiT-XL/2 Peebles & Xie (2023) 1024×4 - 2.27 278.2 675M 250 18
+SiT-XL Ma et al. (2024) 1024×4 - 2.42 238.5 675M 30 2
+SiT-XL w/ Solver Wang et al. (2025) 1024×4 - 2.24 244.1 730M 15 1.2
+AR Taming-VQGAN Esser et al. (2021) 256 1024 15.78 74.3 1.4B 256 8
+RQ-Transformer Huang et al. (2023) 256 16384 7.55 134.0 3.8B 64 8.5
+ViT-VQGAN Yu et al. (2022) 1024 8192 4.17 175.1 1.7B 1024 >10
+LlamaGen-3B Sun et al. (2024) 576 16384 2.18 263.3 3.1B 576 20
+RandAR-XXL Pang et al. (2024) 512 16384 2.15 322.0 1.4B 88 4
+VAR-d30 Tian et al. (2024) 680 4096 1.97 334.7 2.0B 10 0.5
+MVTM MaskGIT Chang et al. (2022) 256 1024 6.18 182.1 227M 8 0.2
+MaskGIL-XXL Xin et al. (2025) 256 16384 3.71 303.4 1.4B 8 0.8
+TiTok-S-128ft. Yu et al. (2024) 128 4096 1.97 281.8 287M 64 1.6
+DDM ITM Hu & Ommer (2024) 1024 16384 5.30 183.0 546M 100 3
+ReDDiT-L (ours) 256 16384 2.13 294.7 346M 20 0.5
+ReDDiT-XL (ours) 256 16384 1.74 313.6 675M 32 1
+ReDDiT-XLf8 (ours) 1024 16384 1.61 318.5 675M 64 2
+Table 2: Comparison of models with the same tokenizer.
+Reconstruction FID (rFID) indicates the tokenizer’s reconstruction quality from its quantized codes.
+Dim denotes codebook dimension.
+AR model’s gFID are indexed from their original report.
+Model VQ Tokenizer Info. Generator
+[…]
+LlamaGen-LAR Sun et al. (2024) LlamaGen-f16 Sun et al. (2024) 2.19 8 343M 3.80
+RandAR-LAR Pang et al. (2024) 343M 2.55
+OursDDM(ReDDiT-L) 346M 2.41
+IBQ-BAR Shi et al. (2025) IBQ-tokenizer Shi et al. (2025) 1.37 256 343M 2.88
+OursDDM(ReDDiT-L) 346M 2.13
+Generation Quality.
+As shown in Tab.
+1, ReDDiT achieves the best performance among the compared discrete models.
+It outperforms the baseline (MaskGIT Chang et al.
+(2022)) with significant margins (gFID: 2.13 vs 6.18 and IS: 294.7 vs.
+182.1).
+It also outperforms the recent DDM method Hu & Ommer (2024) and TiTok-S-128 Yu et al.
+(2024), which is extensively fine-tuned on quantized latents.
+Compared with continuous diffusion models, ReDDiT exhibits on-par efficiency and performance, showing great potential for discrete generation.
+Note that the performance is achieved with a codebook size of 16384, validating ReDDiT’s effectiveness for large-vocabulary codebooks.
+Efficiency.
+ReDDiT is born with the high-efficiency advantage of discrete diffusion models, comparing with AR models.
+As shown in Tab.
+1, the inference time of ReDDiT is slightly longer than MaskGIT, while the performance is overwhelming.
+Without acceleration techniques, ReDDiT achieves a competitive performance which AR and traditional diffusion models use more than 250 steps to achieve.
+Notably, when armed with recent efforts that tailored KV-Cache Liu et al.
+(2025) for discrete diffusion models, ReDDiT’s inference can be further boosted (not included in the main paper for fair comparison).
+See Appendix F for details.
+Besides the major comparison, we also conduct an experiment that utilizes the identical tokenizer in previous AR models and validate our method’s effectiveness.
+As can be seen in Tab.
+2, ReDDiT outperforms AR methods in generation tasks across different tokenizers.
+Note that this comparison is to demonstrate diffusion model’s potential on discretized latents, and current representation alignment methods are inapplicable to AR models due to their unidirectional attention design.
+3.3 Determining Noise Capacity
+Figure 4: Comparison of noise capacities.
+We re-implemented training with MVTM and ReDDiT design with the same training recipe (LlamaGen-f16 as visual tokenizer, with codebook size $16384$).
+The generation quality and representation alignment trends are visualized.
+The reformulated discrete diffusion dynamics defines transitioning from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$.
+Under this setting, it is necessary to empirically determine the optimal value of $m$ for a fixed tokenizer with vocabulary size $d$, as the latent representations learned by VAEs are variant.
+We keep the training setup fixed and conduct experiments w.r.t.
+the noise capacity $m$.
+We also visualize $\mathcal{L}_{\text{RepA}}$, which captures the degree of representation alignment Yu et al.
+(2025) within the transformer.
+The alignment loss visualization shows that increasing the number of absorbing states introduces greater randomness, initially making predictions more difficult due to confusion with valid tokens.
+However, this gap narrows as training progresses, and the model converges to a similar alignment lower bound, suggesting effective representation learning across different configurations.
+As shown in Fig.
+4 (left), generation quality improves with increasing noise capacity initially.
+The LlamaGen-f16 tokenizer achieves peak performance at $m=128$, while the IBQ tokenizer performs best at $m=1024$.
+We attribute this to the codebook design: the lower dimensional LlamaGen-f16 codebook produces more compact latents, which also determines its smaller noise endurance.
+3.4 Ablation Study
+[…]
+Sampling Timeline.
+Figure 5: Illustration of discretized timeline with $K=7$. The slow-to-fast sampling works better than linear schedules.
+Recovering complete information from noise remains critical to diffusion-based models Lu et al.
+(2022); Wu et al.
+(2024).
+Recent work shows MVTM’s non-linear scheduler for training is less critical when using high-capacity tokenizers.
+Evidence of time-invariance in DDMs Sahoo et al.
+(2024); Shi et al.
+(2024) further supports decoupling training from sampling.
+In our experiments, a linear scheduler with constant signal-to-noise ratio decay, yields optimal training dynamics.
+Among the timeline discretization tested, Fig.
+5, the cosine schedule is employed for our ReDDiT model for best performance in Tab.
+3.
+General Design.
+Table 3: Ablated Design choices. ReDDiT-L is trained for 100k iters. Final setting denoted in gray.
+(a) (a) General model design
+Train Config Sample Config gFID Prec. Rec.
+MVTM + RepA loss MVTM sampler 6.83 0.75 0.39
+Switch to objective (11) MVTM sampler 6.23 0.77 0.41
+   same as above Rehash sampler 5.75 0.78 0.45
++ 2D-RoPE + min-SNR Rehash sampler 5.51 0.79 0.45
+   same as above + DFM refine 5.40 0.81 0.52
+(b) (b) Sampling timeline
+Steps Timeline gFID
+20 linear 7.18
+32 linear 6.43
+20 arccos 5.04
+20 square 7.39
+20 cosine 4.91
+We ablate the general choices of ReDDiT, which starts with a re-trained MVTM baseline methods (with LlamaGen-f16 and RepA for faster convergence as default) in Tab.
+3.
+The applied techniques like 2D-RoPE are also ablated with re-training.
+As shown, through the revised objective and our proposed sampler, ReDDiT alone improves FID by $\sim 1.0$ compared to the baseline model.
+When combined with modern modification on transformers, it can further improve the performance, showing its complementaryness with main-stream efforts.
+3.5 Qualitative Result
+Class-conditional Generation.
+Figure 6 presents representative class-conditional samples generated by the proposed ReDDiT model.
+The outputs across diverse image classes consistently exhibit high fidelity and diversity.
+Additional qualitative comparisons and more sample visualizations are provided in Appendix G.
+Image Editing.
+We further demonstrate ReDDiT’s editing capability in Figure 6, highlighting its bi-directional perceptual competence.
+Following MaskGIT Chang et al.
+(2022), we replace a region of the input image with noise tokens and employ the same generation pipeline to inpaint the missing content, conditioned on a class label $c$.
+Thanks to the rehashing noise mechanism, ReDDiT is able to produce diverse and semantically coherent completions without adjusting temperature or other sampling parameters.
+Figure 6: Class-conditional generation and in-painting samples of ReDDiT on ImageNet $256\times 256$.
+4 Conclusion
+We proposed ReDDiT, a discrete visual generative model built upon a discrete diffusion architecture with novel noise designs and efficient sampling strategies.
+Our key contribution lies in the integration of rehashing noise with samplers, which together ensure both diversity and low discrepancy throughout the generative process.
+By introducing rehashing noise, ReDDiT enriches the potential paths that latent variables can traverse during training, regularize training dynamics and enhances model’s representational capacity.
+Extensive experiments demonstrate that discrete generative models can achieve performance on par with their continuous counterparts while offering top-tier efficiency.
+This study paves a promising way for discrete generative modeling and offers fresh insights toward unifying visual and language generation—a path we leave for future exploration.
+Appendix A Use of Large Language Models
+In the process of drafting this paper, Large Language Models (LLMs) were solely utilized for writing polish (e.g., optimizing sentence structure, enhancing expression fluency).
+No LLM was involved in core academic work such as conceptualization, literature review, data analysis, argument construction, or conclusion formulation of this study.
+As the human authors of this paper, we bear full and sole responsibility for the paper’s content, including the accuracy of research data, validity of academic arguments, integrity of research methods, and compliance with academic ethics.
+Appendix B Related Work
+Diffusion Models.
+Diffusion models Ho et al.
+(2020); Song et al.
+(2020) have emerged as a powerful class of generative methods that learn data distributions by reversing a gradual noising process over time.
+These models are primarily designed for continuous domains such as images Dhariwal & Nichol (2021); Gao et al.
+(2023); Peebles & Xie (2023), defining a forward process that transforms data $x_{0}$ into noise $x_{1}$: $x_{t}\sim\mathcal{N}(\sqrt{\alpha_{t}}x_{0};(1-\alpha_{t})\mathbf{I})$ where $\alpha_{t}$ controls the noise schedule.
+The generative (reverse) process learns a denoising model $p_{\theta}(x_{s}\mid x_{t})$, often parameterized via a neural network $\theta$ to predict either noise or clean data.
+Discrete Diffusion Models.
+Discrete diffusion has been previously governed by masked visual token models (MVTMs) Chang et al.
+(2022; 2023); Gu et al.
+(2022); Yu et al.
+(2023; 2024).
+This model leverages a BERT-style [mask] token to corrupt the tokenized image sequence and trained the network with a simple cross-entropy loss on masked tokens, resulting in a score-based prediction.
+It generates tokens in a non-autoregressive fashion, by remasking the tokens with least scores at each inference as depicted in Alg.
+1.
+Recent studies unlocked the principled discrete diffusion model (DDM) Sahoo et al.
+(2024); Shi et al.
+(2024) and discrete flow-matching (DFM) Gat et al.
+(2024); Shaul et al.
+(2024), which adapt the Markov chain theory, enabling generation over text Ou et al.
+(2024); Nie et al.
+(2025), moleculesShaul et al.
+(2024), and other discrete representations Austin et al.
+(2021); Nisonoff et al.
+(2024).
+Unlike MVTMs, the principled DDM and DFM mostly derive a time-weighted cross-entropy loss to supervise the training procedure and apply a gradual unmasking method based on probabilities.
+Appendix C Discrete Diffusion with Rehashing Noise
+Complete Definition and Deduction.
+We provide a full theoretical discussion on the corrupted distribution and reverse process defined in the main paper. The extended discussions with corresponding proofs are marked with teal.
+Given $d$ categories, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be its one-hot vector where the $i$-th value is $1$.
+We denote $\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution, and a basis for absorbing states with capacity $m$: $\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$.
+The sum of $\mathcal{E}$ and $\mathcal{M}$ can be denoted as
+$\mathcal{V}_{(d,m)}\triangleq\left\{\mathbf{v}_{(i,j)}\in\mathbb{R}^{d+m}\,\middle|\,\mathbf{v}_{(i,j)}=\begin{cases}\mathbf{e}_{i}\oplus\mathbf{0}_{m},&\text{for }i=1,\ldots,d,\ j=0\\ \mathbf{0}_{d}\oplus\mathbf{m}_{j},&\text{for }j=1,\ldots,m,\ i=0\end{cases}\right\}.$ (12)
+We further denote the subspace $\mathcal{E}_{d},\ \mathcal{M}_{m}\in\mathcal{V}_{(d,m)}$ which contain either valid or mask tokens, as
+$\mathcal{E}_{d}=\left\{\mathbf{v}_{(i,0)}\in\mathcal{V}_{(d,m)}\,\middle|\,i=1,\ldots,d\right\},\ \mathcal{M}_{m}=\left\{\mathbf{v}_{(0,j)}\in\mathcal{V}_{(d,m)}\,\middle|\,j=1,\ldots,m\right\}.$ (13)
+To exploit visits across all the possible paths, for $0\leq s<t\leq 1$, we write the transition kernel as [^4] To maintain simplicity, we use $\alpha_{t|s}^{\leftarrow}=\frac{\alpha_{t}}{\alpha_{s}}$ and $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$ to denote transition rate for the corruption and reverse process, respectively.
 $q(x_{t}^{i}\mid x_{s}^{i})=\begin{cases}1-\alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=x_{s}^{i},\ x_{s}^{i}\notin\mathcal{M}_{m},\\ 1/m,&\text{if }x_{t}^{i}\in\mathcal{M}_{m},\ x_{s}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise}.\end{cases}$ (14)
-We introduce the essential transition matrix to model the process. Let $\mathbf{I}_{(d,m)}$, $\mathbf{M}_{(d,m)}$ and $\bm{\pi}_{(d,m)}$ be matrices in $\mathbb{R}^{(d+m)\times(d+m)}$, defined as
+Proof of the Corrupted Distribution.
+The presentation in the main paper simplifies the theory without specifying the transition matrix $Q_{t}$ due to page limitation.
+We make a detailed version with important yet basic matrix calculation in this section.
+Let $\mathbf{I}_{(d,m)}$, $\mathbf{M}_{(d,m)}$ and $\bm{\pi}_{(d,m)}$ be matrices in $\mathbb{R}^{(d+m)\times(d+m)}$, defined as
 $\mathbf{I}_{(d,m)}=\begin{bmatrix}I_{d}&0\\ 0&0\end{bmatrix},\quad\mathbf{M}_{(d,m)}=\begin{bmatrix}0&\frac{1}{m}\mathbf{1}_{d\times m}\\ 0&0\end{bmatrix},\quad\bm{\pi}_{(d,m)}=\begin{bmatrix}0&0\\ 0&\frac{1}{m}\mathbf{1}_{m}\mathbf{1}_{m}^{\top}\end{bmatrix}$ (15)
 where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix, and $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of ones.
-The transition matrix $Q_{t|s}\in\mathbb{R}^{(d+m)\times(d+m)}$ corresponding to Eq.
-8 is
+The transition matrix $Q_{t|s}\in\mathbb{R}^{(d+m)\times(d+m)}$ is defined as:
 $Q_{t|s}=\alpha_{t|s}^{\leftarrow}\mathbf{I}_{(d,m)}+(1-\alpha_{t|s}^{\leftarrow})\mathbf{M}_{(d,m)}+\bm{\pi}_{(d,m)}$ (16)
@@
 where $\text{U}(\mathcal{M}_{m}^{L})$ is the uniform distribution on $\mathcal{M}_{m}^{L}$.
-Rehash Sampling.
+Proof of the Reverse Process.
 To generate a sequence of length $L$, the reverse process starts with $x_{1}\sim\text{U}(\mathcal{M}_{m}^{L})$.
-Let $\mathbf{a}\odot\mathbf{b}$ denote the Hadamard product between two vectors $\mathbf{a}$ and $\mathbf{b}$, the reverse process is inferred from D3PM deduction [1] as:
-[…]
+Let $\mathbf{a}\odot\mathbf{b}$ denote the Hadamard product between two vectors $\mathbf{a}$ and $\mathbf{b}$, the reverse process is inferred as:
+$q(x_{s}\mid x_{t})$ $=\dfrac{Q_{t|s}x_{t}\odot Q_{s|0}^{\top}x_{0}}{x_{t}^{\top}Q_{t|0}^{\top}x_{0}}\ \ \ \ \text{(D3PM deduction)}$
 $=\dfrac{[\alpha_{t|s}^{\leftarrow}\mathbf{I}_{(d,m)}x_{t}+(1-\alpha_{t|s}^{\leftarrow})\mathbf{M}_{(d,m)}x_{t}+\bm{\pi}_{(d,m)}x_{t}]\odot[\alpha_{s}x_{0}+(1-\alpha_{s})\mathbf{M}_{(d,m)}^{\top}x_{0}]}{x_{t}^{\top}[\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{M}_{(d,m)}^{\top}x_{0}+\pi_{(d,m)}^{\top}x_{0}]}$
@@
 Case 1.
-For $x_{t}^{i}=x_{0}^{i}$, Eq. 3.2 is simplified as
+For $x_{t}^{i}=x_{0}^{i}$, Eq. C is simplified as
 $q(x_{s}^{i}\mid x_{t}^{i}=x_{0}^{i})$ $=\dfrac{\alpha_{t|s}^{\leftarrow}x_{0}^{i}\odot\alpha_{s}x_{0}^{i}}{\alpha_{t}x_{0}^{i\ \top}x_{0}^{i}}$
@@
 $=\dfrac{(\alpha_{s}-\alpha_{t})\mathbf{M}_{(d,m)}x_{t}^{i}\odot x_{0}+(1-\alpha_{s})\bm{\pi}_{(d,m)}x_{t}^{i}\odot\mathbf{M}_{(d,m)}^{\top}x_{0}}{(1-\alpha_{t})x_{t}^{i\ \top}\mathbf{M}_{(d,m)}^{\top}x_{0}}$ (20)
-Denote $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$ to describe the reverse transition rate, and we have
+Notice that $\alpha_{t|s}^{\rightarrow}=\frac{1-\alpha_{s}}{1-\alpha_{t}}$, and we have
 $q(x_{s}^{i}\in\mathcal{M}_{m}\mid x_{t}^{i}\in\mathcal{M}_{m})=\dfrac{1-\alpha_{s}}{m(1-\alpha_{t})}=\dfrac{\alpha_{t|s}^{\rightarrow}}{m}$ (21)
@@
 Combining case 1 with case 2, we have
-$q(x_{s}^{i}\mid x_{t}^{i})=\begin{cases}1,&\text{if }x_{s}^{i}=x_{t}^{i},\ x_{t}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\rightarrow}/m,&\text{if }x_{s}^{i}\in\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 1-\alpha_{t|s}^{\rightarrow},&\text{if }x_{s}^{i}\notin\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise.}\\ \end{cases}$ (17)
+[…]
 Following MDLM’s deduction, assume that the denoising network can reconstruct $x_{0}$ perfectly, we use $p_{\theta}(x_{t})$ to approximate this reverse process for complex sequences, and get
-$p_{\theta}^{i}(x_{s}^{i}|x_{t})=\begin{cases}1,&\text{if }x_{s}^{i}=x_{t}^{i},\ x_{t}^{i}\notin\mathcal{M}_{m},\\ \alpha_{t|s}^{\rightarrow}/m,&\text{if }x_{s}^{i}\in\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ (1-\alpha_{t|s}^{\rightarrow})p_{\theta}^{i}(x_{t}),&\text{if }x_{s}^{i}\notin\mathcal{M}_{m},\ x_{t}^{i}\in\mathcal{M}_{m},\\ 0,&\text{otherwise.}\\ \end{cases}$ (18)
-The subsequent latents $x_{t}$ are generated by discretizing the reverse timeline $T$ into $K$ steps.
-We denote this schedule as $T^{1:K+1}$ such that $T^{1}=1$ and $T^{K+1}=\varepsilon$, with $\varepsilon$ being an arbitrarily small positive constant.
-Inspired by MD4 [36], we use non-uniform time intervals for efficient sampling.
-Algorithm 1 MVTM Sampling
-1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$,
-2:Settings: codebook entries $\mathcal{V}_{(d,1)}$, number of steps $K$, $G(t)$, Gumbel noise $\mathcal{G}$
-3:Initialize: $x_{1}\leftarrow\mathcal{M}_{1}^{L}$, $t\leftarrow 1$.
-4:for $k=1$ to $K$ do
-5:     $t\leftarrow\frac{K-k+1}{K},s\leftarrow\frac{K-k}{K}$
-6:     $\text{logits}\leftarrow f_{\theta}(x_{t},c)$
-7:     $p_{\text{score}}\leftarrow\text{logits}+G(t)\cdot\mathcal{G}$
-8:     $x_{\text{pred}}\leftarrow\text{argmax}(p_{\text{score}})$triangleright Predict-all
-9:     $x_{s}\leftarrow\text{where}(x_{t}=[\text{m}],x_{\text{pred}},x_{t})$
-10:     $p_{\text{conf}}\leftarrow p_{\text{score}}+G(t)\cdot\mathcal{G}$
-11:     $m_{\text{re}}\leftarrow\text{argsort}(p_{\text{conf}})[1:L\cdot(1-\alpha_{s})]$
-12:     $x_{s}\leftarrow\text{where}(m_{\text{re}},[\text{m}],x_{s})$triangleright Re-mask
-13:end for
-14:Return: fully unmasked sequence $x_{0}$
-Algorithm 2 Rehash Sampling (Ours)
-1:Inputs: label $c$, scheduler $\alpha_{t}$, length $L$.
-2:Settings: codebook entries $\mathcal{V}_{(d,m)}$, number of steps $K$.
-3:Initialize: $x_{1}\leftarrow\text{U}(\mathcal{M}_{m}^{L})$, $t\leftarrow 1$, $T^{1:K}$.
-4:for $k=1$ to $K$ do
-5:     $t\leftarrow T^{k},s\leftarrow T^{k+1}$
-6:    # the rehash operation:
-7:     $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$
-8:     $\text{logits}\leftarrow f_{\theta}(x_{t},c)$
-9:     $p_{0|t}\leftarrow\text{Softmax}(\text{logits})$
-10:     $q_{s|t}\leftarrow\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\cdot p_{0|t}+\delta_{m[0]}\cdot\frac{1-\alpha_{s}}{1-\alpha_{t}}$
-11:     $x_{\text{pred}}\leftarrow\text{Multnm.}(q_{s|t})$triangleright w/ masks
-12:     $x_{s}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},x_{\text{pred}},x_{t})$
-13:end for
-14:Return: fully unmasked sequence $x_{0}$
-In practice, we apply torch.multinomial (Multnm.
-in step 10) for low-discrepancy categorical sampling [18] with the merged probability of absorbing states  [^1] Instead of dividing $\alpha_{t|s}^{\rightarrow}$ and assigning these probabilities to each mask vocabulary.
-We merge the probabilities at step 9 in Alg.
-2 to keep an overall noise sampling probability, as small values might be truncated..
-The random nature of absorbing states inspires a rehash operation: we shuffle these tokens at the beginning of each step by $x_{t}\leftarrow\text{where}(x_{t}\in\mathcal{M}_{m},\text{U}(\mathcal{M}_{m}^{L}),x_{t})$.
-3.3 Discussion
-Comparison with MVTM.
-Masked visual token models (MVTMs) borrow the objective
-$\mathcal{L}_{\text{MVTM}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t}),$ (19)
-from masked language models [7] and predict on masked tokens with a maximum likelihood.
-Besides the reformulated corruption (Eq.
-8) and reverse process (Eq.
-17), ReDDiT differs in the following aspects: (i) the training objective (Eq.
-4), which is derived from the principled DDM which provides better theoretical and empirical results with rehashing noise design.
-(ii) it can easily sample with arbitrarily discretized timeline, while MVTM couples training and inference, restricting its flexibility; (iii) the rehash sampler (Alg.
-2) includes the absorbing states in categorical sampling with lower discrepancy, different from predict-remask sampler with time variant intensity $G(t)$ over Gumbel noise $\mathcal{G}$ (Alg.
-1)  [^2] The logits corresponding to previously restored tokens’ indices are manually set to infinity for both methods, so that they will not be noised again in the following steps.
-This leads to an implementation of any-order auto-regressive model [26] if DDM’s decoded tokens per step is limited to $1$..
-Gumbel-max suffers from numerical inaccuracy [49] and we detect that it worsens on large vocabulary (Fig.
-3 with our reproduced results), which limits the potential of MVTM.
-Relationship to DFM.
-Discrete flow matching (DFM) [11] applies flow matching techniques to discrete structures and introduces a transition process based on masked tokens.
-Its training objective was initially designed as the masked token loss (19), and evolved to a time-weighted cross-entropy loss [34] for generalized diffusion paths, which is similar to ours.
-The similarity enables a direct comparison between the DFM sampler and our rehash sampler using the same trained model weights.
-We noticed that while the DFM sampler offers a refinement mechanism via token-wise updates, it generally requires more steps to reach optimal results.
-However, since the gradual decoding method is shared, we can integrate certain DFM steps into our sampling procedure for refinement.
-We leave this topic for an updated version of our essay.
-Figure 3: Comparison of methods under limited samples.
-Left: Gumbel-max is theoretically equivalent to our method, yet it struggles to reflect the true distribution under limited sample passes.
-The multinomial approach captures the distribution more accurately.
-Right: our model achieves lower gFID across different sampling steps without tuning Gumbel-max, indicating more efficient and faithful sampling.
-(Best viewed in color)
-4 Experiment
-4.1 Implementation
-Datasets.
 […]
-Pre-processing.
-Following the setting in LlamaGen [39], we apply the ten-crop augmentation on images, and use pre-trained tokenizers to convert them to discrete tokens.
-We pick IBQ-f16 [35] tokenizer as default for its scalable and promising performance in generation tasks, which uses a $16\times 16$ downsampling ratio and converts a $256\times 256$ image into 256 discrete tokens.
-The tokenizer has a codebook with 16384 entries.
-The LlamaGen-f16 (used in Tab.
-3) and LlamaGen-f8 tokenizer [39] (used in Tab.
-2) are also used for comparison with recent discrete generation methods.
-All tokenizers are used out-of-the-box without any tuning or modification.
-Training with Representation Alignment.
-Recent study [46] has shown that the alignment of intermediate representations between diffusion transformers and vision encoders accelerates training convergence of diffusion models.
-Accordingly, the alignment is designed as a regularization term with $\lambda=0.5$.
-We extract diffusion transformer’s n-th layer intermediate feature $\textbf{h}^{[n]}(x_{t})$ and align it with the original image’s dinov2-b [25] encoded features $f_{\text{enc}}(x_{0}^{\text{ori}})$.
-The intermediate features are projected by a small trainable MLP $h_{\varphi}$.
-The $\text{sim}(\cdot,\cdot)$ computes the mean of element-wise cosine similarity between embeddings, as
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DDM-linear}}+\lambda\mathcal{L}_{\text{RepA}},\ \ \mathcal{L}_{\text{RepA}}=-\mathbb{E}_{x,\ t}[\ \text{sim}(f_{\text{enc}}(x_{0}^{\text{ori}}),\ h_{\varphi}(\textbf{h}^{[n]}(x_{t})))\ ]\ .$ (20)
-This alignment was proposed for continuous diffusion models, and we firstly validate that it’s also suitable for training discrete models in Tab. 1 (a).
-Model Training.
-The proposed model is based on DiT [28] architecture, with reference to its discrete prediction version [31].
-2D-RoPE [38] and min-SNR [47] are applied for training efficiency.
-The model is optimized using the AdamW optimizer with a cosine decay.
-Training is conducted for 500k iterations on 8 NVIDIA H100 GPUs with a global batch size 1024.
-Class-conditional training is enabled using class embeddings and a drop-rate of 0.1 for generation with classifier-free guidance.
-4.2 Determining Noise Capacity
-Figure 4: Comparison of different noise capacities.
-We re-implemented training with MVTM and ReDDiT design with the same training recipe (LlamaGen-f16 as visual tokenizer, with codebook size $16384$).
-The generation quality and representation alignment trends are visualized.
-The reformulated discrete diffusion dynamics defines transitioning from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$.
-Under this setting, it is necessary to empirically determine the optimal value of $m$ for a fixed tokenizer with vocabulary size $d$, as the latent representations learned by VAEs are variant.
-We keep the training setup fixed and conduct experiments w.r.t.
-the noise capacity $m$.
-We also visualize $\mathcal{L}_{\text{RepA}}$, which captures the degree of representation alignment [46] within the transformer.
-The alignment loss visualization shows that increasing the number of absorbing states introduces greater randomness, initially making predictions more difficult due to confusion with valid tokens.
-However, this gap narrows as training progresses, and the model converges to a similar alignment lower bound, suggesting effective representation learning across different configurations.
-As shown in Fig.
-4 (left), generation quality improves with increasing noise capacity initially.
-The LlamaGen-f16 tokenizer achieves peak performance at $m=128$, while the IBQ tokenizer performs best at $m=1024$.
-We attribute this to the codebook design: the lower dimensional LlamaGen-f16 codebook produces more compact latents, which also determines its smaller noise endurance.
-4.3 Ablation Study
-Table 1: Ablated Design choices. ReDDiT-L is trained for 100k iters. Final setting denoted in gray.
-(a) (a) RepA choices
-Layer RepA gFID
-- $\times$ 7.24
-4 5.72
-8 5.41
-12 5.82
-16 6.12
-(b) (b) General model design
-Configuration gFID Prec. Rec.
-MVTM (19) + RepA 6.83 0.75 0.39
-    objective (20) 6.25 0.76 0.43
-+ 2D-RoPE + min-SNR 5.97 0.78 0.44
-+ rehash sampler 5.53 0.81 0.51
-+ DFM refine 5.40 0.81 0.52
-(c) (c) Sampling timeline
-Steps Timeline gFID
-20 linear 7.18
-32 linear 6.43
-20 arccos 5.04
-20 square 7.39
-20 cosine 4.91
-Unless specified, all the models are trained on ImageNet $256\times 256$ under the default settings for 100k iterations for fair comparison.
-We use a constant guidance scale of 2.0 and 20 steps for generation, and report gFID $\downarrow$ and IS $\uparrow$ computed on 50K samples.
-Precision (Prec.$\uparrow$) and Recall (Rec.$\uparrow$) are also reported in general design for direct diversity comparison.
-General Design.
-We ablate the general choices of ReDDiT, which starts with a re-trained MVTM baseline methods (with LlamaGen-f16 and RepA for faster convergence as default) in Tab.
-1.
-We also tested RepA supervision on different layers of ReDDiT.
-The results are consistent with the original study [46], which indicates that adding RepA objective at the 8-th of 24 layers results in best convergence for L-model.
-The objective and modifications are conducted with re-training.
-We replace the MVTM sampler with the proposed rehash sampler and DFM refinement for further improvement.
-Sampling Timeline.
-Figure 5: Illustration of discretized timeline with $K=7$. The slow-to-fast sampling works better than linear schedules.
-Recovering complete information from noise remains critical to diffusion-based models [22, 42].
-Recent work shows MVTM’s non-linear scheduler for training is less critical when using high-capacity tokenizers.
-Evidence of time-invariance in DDMs [31, 36] further supports decoupling training from sampling.
-In our experiments, a linear scheduler—with constant signal-to-noise ratio decay—yields optimal training dynamics.
-Among the timeline discretization tested, Fig.
-5, the cosine schedule is employed for our ReDDiT model for its optimal performance in Tab.
-1.
-4.4 Performance Comparison
-We compare the proposed ReDDiT model with other generative models on the ImageNet-1K $256\times 256$ in Tab.
-2.
-The IBQ tokenizer is used for the default L and XL models.
-We also utilize LlamaGen-f8 with 128 noise capacity to evaluate its high-resolution potentials (noted as ReDDiT-XLf8).
-We use a linear increasing guidance following the common practice of  [10].
-Table 2: Performance comparison on class-conditional ImageNet 256$\times$256.
-Look-up free quantizers are beyond the scope of this paper.
-ft.
-indicates that the decoder is fine-tuned for quantized latents.
-Wall-clock inference time relative to ReDDiT-XL is reported.
-Type Model Tokenizer Generator
+Appendix D Sampling from Learned Networks
+We present a detailed version of discrete flow matching (DFM) sampler 3 and ours 4, and discuss the integration of them.
+Fig.
+7 presents a quantitative comparison of the vanilla DFM sampler, our proposed rehash sampler, and a hybrid strategy that combines both approaches by incorporating selected DFM steps into the rehash trajectory.
+All methods are evaluated using identical model weights, as the training objectives are compatible due to their shared time-weighted loss formulation.
+The rehash sampler exhibits stronger overall performance than DFM, especially in the 15–32 step range, where it achieves low and stable gFID scores.
+This suggests that our modification enables more efficient decoding trajectories without sacrificing sample quality.
+The hybrid variant, which integrates only the middle and final steps of the DFM update into the rehash schedule, also delivers consistent gains over the vanilla DFM, suggesting that partial refinement from DFM is beneficial even when the majority of the trajectory is governed by our rehash dynamics.
+By leveraging shared gradual decoding infrastructure, the hybrid approach enables practical integration of DFM refinement into the ReDDiT framework with minimal overhead.
+As noted in the main paper, this leads to a $\sim$0.1 improvement in gFID on ImageNet-1K, reinforcing the complementary strengths of the two samplers.
+We leave the comprehensive study on the optimal integration of different samplers for future exploration.
+Figure 7: Generation quality comparison with DFM methods. The experiments are conducted on ReDDiT-L with a constant classifier-free guidance ($\text{cfg}=2.0$).
+Algorithm 3 DFM Sampling Stepwise Pseudo Code
 […]
-Diff. LDM-4 [29] 4096×3 - 3.60 247.7 400M 250
-DiT-XL/2 [28] 1024×4 - 2.27 278.2 675M 250 18
-MDTv2 [10] 1024×4 - 1.58 314.7 676M 250 >18
-MAR-H [20] 256×16 - 1.55 303.7 943M 256 10
-MVTM MaskGIT [4] 256 1024 6.18 182.1 227M 8 0.2
-TiTok-S-128ft. [45] 128 4096 1.97 281.8 287M 64 1.6
-AR Taming-VQGAN [9] 256 1024 15.78 74.3 1.4B 256 8
-RQ-Transformer [17] 256 16384 7.55 134.0 3.8B 64 8.5
-ViT-VQGAN [43] 1024 8192 4.17 175.1 1.7B 1024 >10
-LlamaGen-3B [39] 576 16384 2.18 263.3 3.1B 576 20
-RandAR-XXL [27] 512 16384 2.15 322.0 1.4B 88 4
-VAR-d30 [41] 680 4096 1.97 334.7 2.0B 10 0.5
-DDM ITM [16] 1024 16384 5.30 183.0 546M 100 3
-ReDDiT-L (ours) 256 16384 2.13 294.7 346M 20 0.5
-ReDDiT-XL (ours) 256 16384 1.74 313.6 675M 32 1
-ReDDiT-XLf8 (ours) 1024 16384 1.61 318.5 675M 64 2
-Table 3: Comparison of models with the same tokenizer.
-Reconstruction FID (rFID) indicates the tokenizer’s reconstruction quality from its quantized codes.
-Dim denotes codebook dimension.
-AR model’s gFID are indexed from their original report.
-Model VQ Tokenizer Info. Generator
+2:Compute jump probabilities: $j_{t}\leftarrow 1-\alpha_{t}$, $j_{s}\leftarrow 1-\alpha_{t-\Delta t}$
+3:Determine guidance scale $\omega$ from schedule
+4:Obtain logits $\text{logits}_{\text{cond}}$, $\text{logits}_{\text{uncond}}$ via forward pass
 […]
-LlamaGen-LAR [39] LlamaGen-f16 [39] 2.19 8 343M 3.80
-RandAR-LAR [27] 343M 2.55
-OursDDM(ReDDiT-L) 346M 2.41
-IBQ-BAR [35] IBQ-tokenizer [35] 1.37 256 343M 2.88
-OursDDM(ReDDiT-L) 346M 2.13
-Generation Quality.
+[…]
+7:Sample $\hat{x}_{0}\sim p_{x_{0}}$ using categorical sampling
+8:Construct one-hot encodings: $\delta_{x_{0}},\delta_{x_{t}}$
+[…]
+[…]
+11:Overwrite $u$ in masked range with corrective terms
+12:Mask entries already present in $x_{t}$ from $u$
+13:Compute total transition intensity: $\lambda\leftarrow\sum u$, elementwise
+14:Draw Bernoulli mask: $M\sim\text{Bernoulli}(1-\exp(-\lambda))$
+15:For each masked position in $M$, sample from categorical $u$ to obtain updated $x_{s}$
+16:return $x_{s}$
+Algorithm 4 Rehash Sampling (ours) Stepwise Pseudo Code
+[…]
+2:Indentify the masked tokens: $M\leftarrow[x_{t}\in\mathcal{M}_{m}]$
+3:Rehash $x_{t}$: $x_{t}\leftarrow M\cdot\texttt{random\_shuffle}(\text{mask\_vocab})+(1-M)\cdot x_{t}$.
+4:Compute move coefficients: $k_{t}\leftarrow 1-\alpha_{t}$, $k_{s}\leftarrow 1-\alpha_{t-\Delta t}$
+5:Determine guidance scale $\omega$ from schedule
+6:Obtain logits $\text{logits}_{\text{cond}}$, $\text{logits}_{\text{uncond}}$ via forward pass
+[…]
+[…]
+9:Set mask probability: $p_{\text{mask}}\leftarrow k_{s}$
+10:Construct proposal distribution: $q_{x_{s}}\leftarrow\frac{k_{t}-k_{s}}{k_{t}}\cdot p_{x_{0}}$
+11:Overwrite mask token logits:
+$q_{x_{s}}[:,:,m{:}]\leftarrow 0,\quad q_{x_{s}}[:,:,m]\leftarrow\frac{p_{\text{mask}}}{k_{t}}$
+where $m$ is the start of mask token index
+12:Sample $\hat{x}\sim q_{x_{s}}$ using categorical sampling
+13:Identify preserved tokens: $c\leftarrow[x_{t}<m]$
+14:Combine result: $x_{s}\leftarrow c\cdot x_{t}+(1-c)\cdot\hat{x}$
+15:return $x_{s}$
+Appendix E Experiment Details
+We provide detailed training and generation configurations for ReDDiT in Table 4.
+Our method incorporates DINOv2-B for representation alignment, which requires computing image features during the forward pass (only activated during training).
+This introduces an overhead, making training roughly 1.2× slower than solely on discrete tokens.
+However, this additional cost is offset by faster convergence and improved stability, particularly in early training stages.
+The use of quantized latents allows for larger batch sizes under limited GPU memory, making our approach more accessible for low-resource settings.
+Additionally, aligning discrete codes with semantic features improves the quality and diversity of learned representations.
+Overall, our design balances computational efficiency with model performance, making it a practical choice for both research and deployment.
+Table 4: Experiment details for ReDDiT on ImageNet-1K. Vari. refers to a time-variant growing guidance scale following MDTv2, which is a common practice for diffusion models.
+Setting ReDDiT-L (Ablation) ReDDiT-L ReDDiT-XL ReDDiT-XLf8
+Hidden Size 1024 1024 1280 1280
+Transformer Block 24 24 28 28
+Attention Head 16 16 20 20
+Image Tokenizer LlamaGen-f16 IBQ-f16 IBQ-f16 LlamaGen-f8
+Codebook Size 16384 16384 16384 16384
+Noise Capacity 128 1024 1024 128
+Sequence Length 256 256 256 1024
+RepA Latent Size 16×16 16×16 16×16 32×32
+Batch Size 64 64 32 16
+Global Batch Size 1024 1024 1024 1024
+LR scheduler Cosine Decay Cosine Decay Cosine Decay Cosine Decay
+Learning Rate 3e-4 3e-4 3e-4 4e-4
+Minimal LR 1e-5 1e-5 1e-5 1e-5
+Warmup Steps 2k 2k 2k 2k
+Training Steps 500k 500k 500k 500k
+Training Time $\sim$1 day $\sim$1 day $\sim$2 days $\sim$3 days
+Generation CFG (Vari.) 1.0-5.0 1.0-6.5 1.0-6.5 1.0-5.5
+Appendix F Accelerating ReDDiT
+Recent efforts on scaling and accelerating discrete diffusion models are making this generative paradigm more practical than theoretical attempts.
+We adapt the dLLM-Cache Liu et al.
+(2025) design into our framework, which efficiently reuses intermediate computations without compromising model performance.
+Since the condition is modulated using AdaLN and introduces minimal calculation, we do not activate $K_{p}$ (cache for prompt).
+As the decoding of visual sequence varies with time more quickly than language decoding, we implement the cache for response with small values like $K_{r}=2\ \text{or}\ 4$, which means the $K$ and $V$ of transformer layer is updated every $2$ or $4$ decoding steps instead of per step.
 As shown in Tab.
-2, ReDDiT achieves the best performance among the compared discrete models.
-It outperforms the baseline (MaskGIT [4]) with significant margins (gFID: 2.13 vs 6.18 and IS: 294.7 vs.
-182.1).
-It also outperforms the recent DDM method [16] and TiTok-S-128 [45], which is extensively fine-tuned on quantized latents.
-Compared with continuous diffusion models, ReDDiT exhibits higher efficiency and on-par performance, showing great potential for discrete generation.
-Note that the performance is achieved with a codebook size of 16384, validating ReDDiT’s effectiveness for large-vocabulary codebooks.
-Efficiency.
-ReDDiT is born with the high-efficiency advantage of discrete diffusion models.
-As shown in Tab.
-2, the inference time of ReDDiT is slightly longer than MaskGIT, while the performance is overwhelming.
-With only 20 steps it achieves a competitive performance which diffusion models uses more than 250 steps to achieve.
-It also demonstrates comparable results with significantly smaller model size than recent VAR [41] method.
-Besides the major comparison, we also conduct an experiment that utilizes the identical tokenizer in previous auto-regressive (AR) works and validate our method’s effectiveness.
-As can be seen in Tab.
-3, ReDDiT outperforms AR methods in generation tasks across different tokenizers.
-4.5 Qualitative Results
-Figure 6 presents representative class-conditional samples generated by the proposed ReDDiT model.
-The outputs across diverse image classes consistently exhibit both high fidelity and diversity.
-We further demonstrate ReDDiT’s editing capability in Figure 7, highlighting its bi-directional perceptual competence.
-Following MaskGIT [4], we replace a region of the input image with noise tokens and employ the same generation pipeline to inpaint the missing content, conditioned on a class label $c$.
+5, the inference speed is boosted up to 2 times with minimal performance drop, which makes our largest model ReDDiT-XLf8 comparable to diffusion models with accelerated solvers.
+Table 5: Acceleration of ReDDiT using response cache $K_{r}$.
+Model Config Performance
+[…]
+ReDDiT-L 32 2 $\times 1.33$ 2.28 ($\Delta=0.15$)
+ReDDiT-XL $\times 1.52$ 1.88 ($\Delta=0.14$)
+ReDDiT-XL 64 4 $\times 2.17$ 1.83 ($\Delta=0.09$)
+ReDDiT-XLf8 $\times 2.56$ 1.71 ($\Delta=0.10$)
+Appendix G Qualitative Results
+We provide a comparison between MVTM and our method on generated images, and more samples of ReDDiT’s generation in Fig. 8.
 Figure 8: Upper: Comparison between MVTM and our method on generated images. Below: Class-conditional generation samples of ReDDiT on ImageNet $256\times 256$.
-Figure 7: Class-conditional generation and in-painting samples of ReDDiT on ImageNet $256\times 256$.
-5 Conclusion
-We proposed ReDDiT, a discrete visual generative model built upon a principled discrete diffusion architecture with novel noise designs and efficient sampling strategies.
-Our key contribution lies in the integration of rehashing noise and samplers, which together ensure both diversity and low discrepancy throughout the generative process.
-By introducing rehashing noise, ReDDiT enriches the potential paths that latent variables can traverse during training, enhancing representational capacity.
-Extensive experiments demonstrate that discrete generative models can achieve performance on par with their continuous counterparts while offering significantly improved efficiency.
-This work paves a promising way for discrete generative modeling and offers fresh insights toward unifying visual and language generation—a path we leave for future exploration.